### v1

# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

Abstract
In this paper, we propose a new class of positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra. We focus on kernels whose inputs and outputs are functions and generalize existing kernels, such as polynomial, product, and separable kernels, by introducing a truncation parameter $n$ that describes the noncommutativity of the products appearing in the kernels. When $n$ goes to infinity, the proposed kernels tend to the existing commutative kernels. If $n$ is finite, they exhibit different behavior, and the noncommutativity induces interactions along the data function domain. We show that the truncation parameter $n$ is a governing factor leading to performance enhancement: by setting an appropriate $n$, we can balance the representation power and the complexity of the representation space. The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels.
1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning [1, 2, 3, 4, 5]. They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by function spaces called reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered [1, 6].
Standard positive definite kernels are real- or complex-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector-, function-, or operator-valued outputs have also been investigated [4, 7, 8]. The kernel, in this case, is instead an operator- or function-valued positive definite function, and the associated feature space is a vector-valued RKHS (vvRKHS) [9, 7, 10] or a reproducing kernel Hilbert $C^{*}$-module (RKHM) [8]. RKHM is a generalization of RKHS by means of $C^{*}$-algebra, and it also generalizes vvRKHS. $C^{*}$-algebra is a generalization of the space of complex values and has a structure of the norm, product, and involution [11, 12]. It unifies operators and functions. Thus, the operator- and function-valued kernels for vvRKHSs and RKHMs can be unified as $C^{*}$-algebra-valued kernels. It has been shown that the norms in $C^{*}$-algebras alleviate the dependency of the generalization bound on the output dimension [13] and the product structure of a $C^{*}$-algebra goes beyond the convolution operation [14].
One big challenge for kernel methods is the choice of positive definite kernels. Especially for vvRKHSs and RKHMs, the construction of good $C^{*}$-algebra-valued kernels is not obvious. A typical kernel is the separable kernel, which is defined by the product of a real- or complex-valued kernel and a positive semi-definite operator [4]. Attempts have been made to construct positive definite kernels that go beyond separable kernels. Huusari and Kadri [15] proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. [14] proposed to use the product of circulant matrices and general squared matrices to construct positive definite kernels. By using the kernel, we can generalize the convolution and capture the effect of interactions of different Fourier components.
In this paper, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the field of noncommutative geometry and $C^{*}$-algebra [16, 17, 18]. We focus on the case of functional inputs and outputs, and consider the $C^{*}$-algebra of continuous functions. We generalize typical existing kernels: polynomial, product, and separable kernels using the pair of maps called the approximate order isomorphism, which takes an important role in the spectral truncation [19, 17, 20]. The proposed kernels are parameterized by a natural number $n$ that corresponds to the dimension of the truncated space. We approximate functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices. Thus, $n$ describes the resolution of the discretization of the function, and we call it the truncation parameter. We show that the proposed kernels converge to the existing kernels as $n$ goes to infinity.
To analyze functional data, we have been focused on how to interpolate observed data to obtain functions [21, 22, 8]. On the other hand, our results shed light on good effects of the discretization (setting $n$ as a finite number) on learning process. Indded, the truncation parameter $n$ takes an important role from at least two perspectives. First, one of the important structures in the existing kernels is the product, and $n$ describes the noncommutativity of the product. Although the product of two functions is too simple to capture structures of these functions, the product of two approximated Toeplitz matrices is more complex, which enables us to capture the structures of the functions. More precisely, for two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$. Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product. On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$. This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$. Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHM. We show that if $n$ is small, then the representation power is low, and the model complexity is small. On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large. The expectation loss for learning a function is decomposed into the empirical loss, controlled by the representation power, and the complexity. By introducing the parameter $n$ and setting $n$ to balance the empirical loss and the complexity, we obtain higher performance compared to the case of existing kernels.
Our contributions are summarized as follows:
  • We propose a new class of function-valued positive definite kernels that generalize existing kernels: polynomial, product, and separable kernels. The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter $n$, which allows us to introduce noncommutativity into the learning process. The proposed kernels induce the interactions along the data function domain. (Definition 6 and Subsection 3.1)
  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels. The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHMs. (Section 4)
  • We apply the proposed kernels to kernel ridge regression. The result illustrates the advantage of introducing the truncation parameter $n$ in practical settings. (Section 5)
We emphasize that although we focus on the case where the inputs and outputs are functions and three types of kernels, our idea of constructing new positive definite kernels is general. We establish a new framework that allows us to go beyond the usual commutative kernels and to assign a significance to noncommutative structures by discretizing functions.
2 Preliminaries
We review mathematical notions required in this paper.
2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
$C^{*}$-algebra is a Banach space equipped with a product and an involution satisfying the $C^{*}$ identity (condition 3 below). It is a natural generalization of the space of complex numbers.
Definition 1 ($C^{*}$-algebra)
A set $\mathcal{A}$ is called a $C^{*}$-algebra if it satisfies the following conditions:
  1. 1.
    $\mathcal{A}$ is an algebra over $\mathbb{C}$ and equipped with a bijection $(\cdot)^{*}:\mathcal{A}\to\mathcal{A}$ that satisfies the following conditions for $\alpha,\beta\in\mathbb{C}$ and $a,b\in\mathcal{A}$:
    $\bullet$ $(\alpha a+\beta b)^{*}=\overline{\alpha}a^{*}+\overline{\beta}b^{*}$,  $\bullet$ $(ab)^{*}=b^{*}a^{*}$,  $\bullet$ $(a^{*})^{*}=a$.
  2. 2.
    $\mathcal{A}$ is a normed space endowed with $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds. In addition, $\mathcal{A}$ is complete with respect to $\|\cdot\|_{\mathcal{A}}$.
  3. 3.
    For $a\in\mathcal{A}$, the $C^{*}$ identity $\|a^{*}a\|_{\mathcal{A}}=\|a\|_{\mathcal{A}}^{2}$ holds.
If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit, and denoted by $1_{\mathcal{A}}$.
In this paper, we focus on the $C^{*}$-algebra of continuous functions.
Example 1
Let $\mathbb{T}=\mathbb{R}/2\pi\mathbb{Z}$ be the torus and $C(\mathbb{T})$ be the space of continuous functions on $\mathbb{T}$. Then, $\mathcal{A}:=C(\mathbb{T})$ is a $C^{*}$-algebra by means of the product: $(cd)(z)=c(z)d(z)$, involution: $c^{*}(z)=\overline{c(z)}$, and norm: $\|c\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}}|c(z)|$ for $c,d\in\mathcal{A}$.
We review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
Definition 2 (Positive)
An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds. For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive. We denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
Definition 3 (Infimum and minimum)
For a subset $\mathcal{S}$ of $\mathcal{A}$, $a\in\mathcal{A}$ is said to be a lower bound with respect to the order $\leq_{\mathcal{A}}$, if $a\leq_{\mathcal{A}}b$ for any $b\in\mathcal{S}$. Then, a lower bound $c\in\mathcal{A}$ is said to be an infimum of $\mathcal{S}$, if $a\leq_{\mathcal{A}}c$ for any lower bound $a$ of $\mathcal{S}$. If $c\in\mathcal{S}$, then $c$ is said to be a minimum of $\mathcal{S}$.
We now define RKHM. Let $\mathcal{X}$ be a non-empty set for data. To construct an RKHM, we first introduce $\mathcal{A}$-valued positive definite kernel.
Definition 4 ($\mathcal{A}$-valued positive definite kernel)
An $\mathcal{A}$-valued map $k:\mathcal{X}\times\mathcal{X}\to\mathcal{A}$ is called a positive definite kernel if it satisfies the following conditions:
$\bullet$ $k(x,y)=k(y,x)^{*}$  for $x,y\in\mathcal{X}$,
$\bullet$ $\sum_{i,j=1}^{n}c_{i}^{*}k(x_{i},x_{j})c_{j}$ is positive semi-definite  for $n\in\mathbb{N}$, $c_{i}\in\mathcal{A}$, $x_{i}\in\mathcal{X}$.
Let $\phi:\mathcal{X}\to\mathcal{A}^{\mathcal{X}}$ be the feature map associated with $k$, defined as $\phi(x)=k(\cdot,x)$ for $x\in\mathcal{X}$ and let $\mathcal{M}_{k,0}=\{\sum_{i=1}^{n}\phi(x_{i})c_{i}|\ n\in\mathbb{N},\ c_{i}\in\mathcal{A},\ x_{i}\in\mathcal{X}\ (i=1,\ldots,n)\}$. We can define an $\mathcal{A}$-valued map $\big{\langle}\cdot,\cdot\big{\rangle}_{k}:\mathcal{M}_{k,0}\times\mathcal{M}_{k,0}\to\mathcal{A}$ as
$\bigg{\langle}\sum_{i=1}^{n}\phi(x_{i})c_{i},\sum_{j=1}^{l}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{n}\sum_{j=1}^{l}c_{i}^{*}k(x_{i},y_{j})d_{j},$
which enjoys the reproducing property $\left\langle\phi(x),f\right\rangle_{k}=f(x)$ for $f\in\mathcal{M}_{k,0}$ and $x\in\mathcal{X}$.
The norm in $\mathcal{M}_{k,0}$ is defined as $\|f\|_{k}^{2}=\|\left\langle f,f\right\rangle_{k}\|_{\mathcal{A}}$, and we can also define the $\mathcal{A}$-valued absolute value $|\cdot|_{k}$ as $|f|_{k}^{2}=\left\langle f,f\right\rangle_{k}$. The reproducing kernel Hilbert $\mathcal{A}$-module (RKHM) $\mathcal{M}_{k}$ associated with $k$ is defined as the completion of $\mathcal{M}_{k,0}$ with respect to $\|\cdot\|_{k}$. See, for example, the references [12, 11, 23, 8] for more details about $C^{*}$-algebra and RKHM.
2.2 Spectral truncation on the torus
The product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure. Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
$R_{n}(x)_{j,l}=\left\langle e_{j},M_{x}e_{l}\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}}x(t)\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{d}t.$
Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix [24]. It is characterized only by a vector $r_{n}(x)\in\mathbb{C}^{2n-1}$, where the $(j-l)$th element of $r_{n}(x)$ is $R_{n}(x)_{j,l}$. Note that $R_{n}(x)_{j,l}$ is regarded as the $(j-l)$th Fourier component of $x$, and the vector $r_{n}(x)$ is the coordinate of the function $x$ projected on the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$.
For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=1/n\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$. The map $S_{n}$ takes the representation matrix $R_{n}(x)$ back to a function that approximates the original function $x$. Indeed, we have
$S_{n}(R_{n}(x))(z)=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j,l=0}^{n-1}\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{e}^{\mathrm{i}(j-l)z}\mathrm{d}t=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}l(z-t)}\mathrm{d}t=x\ast F_{n}(z),$
where $F_{n}(t)=1/n\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ [25, 17].
Proposition 5
For each $z\in\mathbb{T}$, $x\ast F_{n}(z)\to x(z)$ as $n\to\infty$.
The pair of maps $R_{n}$ and $S_{n}$ is called the approximate order isomorphism.
3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. The following three kinds of $\mathcal{A}$-valued kernels are often considered in existing studies [14, 13, 1, 9, 4].
Example 2
For $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$,
  1. 1.
    polynomial kernel $k^{\operatorname{poly},q}(x,y)(z)=\sum_{i=1}^{d}\alpha_{i}(\overline{x_{i}(z)}y_{i}(z))^{q}$,
  2. 2.
    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$,
  3. 3.
    separable kernel $k^{\operatorname{sep},q}(x,y)=\tilde{k}(x,y)\prod_{j=1}^{q}(a_{j}^{*})^{q}a_{j}^{q}$.
Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ and $\tilde{k}:\mathcal{A}^{d}\times\mathcal{A}^{d}\to\mathbb{C}$ are complex-valued continuous positive definite kernels. In addition, $a_{j}\in\mathcal{A}$ is the parameter of $k^{\operatorname{sep},q}$.
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the existing kernels. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
Definition 6
With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, let
$k_{n}^{\operatorname{poly},q}(x,y)$ $=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)},$
$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)},$
$k_{n}^{\operatorname{sep},q}(x,y)$ $=\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j})\bigg{)}.$
Here, $A^{*}$ for a matrix $A$ is the adjoint, and we denote by $\tilde{k}_{i,j}(x,y)$ the map $z\mapsto\tilde{k}_{i,j}(x(z),y(z))$.
In Definition 6, to construct the kernels, we first project functions in $\mathcal{A}$ onto the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ using the map $R_{n}$ and obtain Toeplitz matrices. Then, we consider the product of the Toeplitz matrices. Note that the product of two Toeplitz matrices is not always a Toeplitz matrix. Finally, we apply $S_{n}$ to the matrix and take it back to the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ and obtain the output function of the kernel. Figure 1 schematically shows the construction of the simplest kernel $k_{n}^{\operatorname{poly},1}(x,y)=S_{n}(R_{n}(x)^{*}R_{n}(y))$ defined in Definition 6.
Figure 1: Overview of the construction of the simplest kernel $k_{n}^{\operatorname{poly},1}(x,y)=S_{n}(R_{n}(x)^{*}R_{n}(y))$
By the definitions of $S_{n}$ and $R_{n}$, we can show the following identity for $x_{1},\ldots,x_{q},y_{1},\ldots,y_{q}\in\mathcal{A}$ (see the proof of Theorem 7 in Appendix A):
$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (1)
[…]
$F_{n}^{q,P}(t)=1/n\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}$ (2)
for $t\in\mathbb{T}^{q}$ is the Fejér kernel on $\mathbb{T}^{q}$ with respect to the polyhedron $P$. Here, for $j\in\mathbb{Z}$, $jP=\{jz\,\mid\,z\in P\}$. The generalization of the Fejér kernel on $\mathbb{T}$, which is defined in Subsection 2.2, to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated [26, 25]. Note that $F_{n}^{q,P}(t)$ is a real-valued function since for $r\in P$, we have $-r\in P$. Thus, if a kernel in Example 2 is real-valued, then the corresponding kernel in Definition 6 is also real-valued. Eq. (1) implies that the value of $S_{n}\big{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y_{j})\big{)}$ at $z$ is determined not only by $x_{1}(z),\ldots,x_{q}(z),y_{1}(z),\ldots,y_{q}(z)$, but also by $x_{1}(t_{1}),\ldots,x_{q}(t_{q}),y_{1}(t_{q+1}),\ldots,y_{q}(t_{2q})$ with different values $t_{1},\ldots,t_{2q}$ from $z$.
3.1 Convergence and interactions
We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2, respectively. In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of existing kernels. In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$. In the following, all the proofs of theoretical statements are documented in Appendix A.
Theorem 7
For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
As we can see in Eq. (1), if the kernel involves $q$ products of Toeplitz matrices generated by $R_{n}$, it is represented by a Fejér kernel on $\mathbb{T}^{q}$. To show Theorem 7, we apply the following lemmas.
Lemma 8
[…]
Lemma 9
Let $P$ be a convex polyhedron. Let $F_{n}^{q,P}$ be the Fejér kernel on $\mathbb{T}^{q}$ defined as Eq. (2). Then, for any $g\in C(\mathbb{T}^{q})$ and any $z\in\mathbb{T}^{q}$, $g\ast F_{n}^{q,P}(z)\to g(z)$ as $n\to\infty$.
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 8 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (1). Lemma 9 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini and Travaglini [25]. We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (1), and apply Lemmas 8 and 9 to show Theorem 7. See Appendix A for more details.
As seen in Theorem 7, the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows. On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel. Regarding the relationship between $q$ and the interactions, we have the following remark.
Remark 10
We expect that the convergence in Theorem 7 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger. This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$. See Appendix B for more details.
3.2 Positive definiteness
To define an RKHM introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product. Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
Proposition 11
The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
As for $k_{n}^{\operatorname{prod},q}(x,y)$, $x$ and $y$ depend on $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products. Thus, we cannot show its positive definiteness. However, we can modify it to become positive definite as follows.
Proposition 12
Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
$\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
To set the value of the parameter $\beta_{n}$, we have the following bound.
Lemma 13
We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
Remark 14
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, the bound in Lemma 13 may not be tight, and even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated [27, 28]. Indeed, in practical computations in Subsection 5.1, the algorithms can work even if we set $\beta_{n}<n^{q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. To derive the generalization bound, we derive an upper bound of the Rademacher complexity. Let $\Omega$ be a probability space with the probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be input and output samples from a distributions of $\mathcal{A}_{0}^{d}$-valued and $\mathcal{A}_{1}$-valued random variables $X$ and $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d. Rademacher variables, which take their values on $\{-1,1\}$. For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$. Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$. An $\mathcal{A}$-valued version of the Rademacher complexity is also defined by Hashimoto et al. [14] as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$. Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3.
Theorem 15
Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$ (3)
for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
The first term of the bound (3) is the empirical loss, and the second term of the bound represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$ associated with $k_{n}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$ in $\mathcal{A}$ is larger, $\mathcal{M}_{k_{n}}$ can represent more functions. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, according to Theorem 15, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger. An advantage of introducing the parameter $n$ and constructing the new kernels based on the existing kernels in Example 2 is that we can control the empirical loss and the complexity through $n$. By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$ small.
To show Theorem 15, we use the $\mathcal{A}$-valued Rademacher complexity. Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity.
Proposition 16
Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
5 Application to Kernel Ridge Regression
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$. Consider the typical minimization problem for regression, which is also considered in [14]:
$\min_{f\in\mathcal{M}_{k}}\bigg{(}\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{\mathcal{A}}^{2}+\lambda|f|_{k}^{2}\bigg{)},$ (4)
where $\lambda\geq 0$ is the regularization parameter. We apply the approximation version of the representer theorem for RKHMs over general $C^{*}$-algebras (see Appendix C). Then, the solution of the problem (4) is approximated by a vector having the form of $\sum_{i=1}^{N}\phi_{i}(x_{i})c_{i}$. The problem is reduced to
$\min_{c_{j}\in\mathcal{A}}\bigg{(}\sum_{i=1}^{N}\bigg{|}\sum_{j=1}^{N}k(x_{i},x_{j})c_{j}-y_{i}\bigg{|}_{{\mathcal{A}}}^{2}+\lambda\bigg{|}\sum_{j=1}^{N}\phi(x_{j})c_{j}\bigg{|}_{k}^{2}\bigg{)}$
$\quad=\min_{c_{j}\in\mathcal{A}}(\mathbf{c}^{*}\mathbf{G}^{2}\mathbf{c}-\mathbf{c}^{*}\mathbf{G}\mathbf{y}-\mathbf{y}^{*}\mathbf{G}\mathbf{c}+\lambda\mathbf{c}^{*}\mathbf{G}\mathbf{c}),$ (5)
where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$. Then, the solution $\mathbf{c}$ of the problem (5) satisfies $\mathbf{y}=(\mathbf{G}+\lambda I)\mathbf{c}$, which means $\mathbf{y}(z)=(\mathbf{G}(z)+\lambda I)\mathbf{c}(z)$. Therefore, for each $z\in\mathbb{T}$, we obtain $\mathbf{c}(z)$ by computing $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$.
Computational cost
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. On the other hand, that for $k^{\operatorname{poly},q}$, $\hat{k}^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ is $O(qmN^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use a vvRKHS. For vvRKHSs, to obtain $m$ values as an output, we generally apply a $\mathbb{C}^{m\times m}$-valued kernel. In this case, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$. We can reduce the cost with respect to $N$, e.g., by Nyström method [29]. Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
5.1 Numerical results
We show numerical results that illustrate the effects of the noncommutativity on the learning process.
5.1.1 Experiment with synthetic data
To observe the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compare them to the existing kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$), we consider a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ are sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1/\Delta^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=\mathrm{e}^{\sin z}$. We estimated $f$ using kernel ridge regression. Figure 2 (a-c) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the empirical error and the complexity in Eq. 3 are balanced, and the best possible expected error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the existing kernel ($n=\infty$). We have similar results for the other two kernels. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 14, $\hat{k}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. See Appendix D.1 for more details.
(a) $k={k}_{n}^{\operatorname{poly},q}$
(b) $k=\hat{k}_{n}^{\operatorname{prod},q}$
(c) $k=k_{n}^{\operatorname{sep},q}$
(d) $k=\hat{k}_{n}^{\operatorname{prod},q}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
(e) Output images
Figure 2: (a-c) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the image recovering task with different values of $n$. (Box plot of results of five independent runs with different training and test data.) (e) Output images of the image recovering task with different values of $n$.
5.1.2 Experiment with MNIST
We consider an image recovering task with MNIST [30]. For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We try to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. Figure 2 (d) and (e) show the test error and the output images with different values of $n$. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. As for the output images, when $n=\infty$, i.e., the existing kernel $k^{\operatorname{prod},q}$, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
6 Generalization to Other $C^{*}$-algebras
Setting approximate order isomorphisms for more general $C^{*}$-algebras has been investigated. Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
Continuous functions on high-dimensional torus [20]
[…]
Continuous functions on the sphere [19]
For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space. Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$. We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
7 Conclusion and Limitation
In this paper, we proposed a new class of positive definite kernels for functional inputs and outputs based on the spectral truncation. The proposed kernels are indexed by a truncation parameter $n$, and they converge to existing kernels: polynomial, product, and separable kernels, as $n$ goes to infinity. Setting $n$ as a finite value, we can introduce noncommutativity into the learning process. More precisely, we can induce interactions along the data function domain and control the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. Since $n$ describes the resolution of the discretization functions, our results show the positive effect of discretization on generalization. We applied the proposed kernel to kernel ridge regression and numerically observed that the proposed kernels outperformed the existing kernels.
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although we can generalize this setting and define positive definite kernels for more general inputs and outputs as we discussed in Section 6, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
Acknowledgements
The representer theorem stated in Appendix C is pointed out by Ms. Miho Mukohara. We would like to thank her for the helpful discussion. We would also like to thank Mr. Noboru Isobe for a constructive discussion. Hachem Kadri is partially supported by grant ANR-19-CE23-0011 from the French National Research Agency. Masahiro Ikeda is partially supported by grant JPMJCR1913 from JST CREST.
Appendix
Appendix A Proofs
We provide the proofs of theorems, propositions, and lemmas stated in the main text.
Theorem 7    For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)\to k^{\operatorname{poly},q}(x,y)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are the same. We have
$k_{n}^{\operatorname{poly},q}(x,y)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q}=0}^{n-1}R_{n}(x_{i})^{*}_{j,r_{1}}\cdots R_{n}(x_{i})^{*}_{r_{q-1},r_{q}}R_{n}(y_{i})_{r_{q},r_{q+1}}\cdots R_{n}(y_{i})_{r_{2q-1},l}\mathrm{e}^{\mathrm{i}(j-l)z}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q-1}=0}^{n-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)t_{1}}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})t_{q}}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})t_{q+1}}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})t_{2q}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}\mathrm{e}^{\mathrm{i}(j-l)z}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q-1}=0}^{n-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=0}^{n-1}\bigg{(}\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{m}-\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{m-1}\bigg{)}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=0}^{n-1}\sum_{\begin{subarray}{c}j\lor l\lor r_{1}\lor\ldots\lor r_{2q-1}=m\\ 0\leq j,l,r_{1},\ldots,r_{2q-1}\leq m\end{subarray}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}r_{1}(z-t_{1})}\cdots\mathrm{e}^{\mathrm{i}r_{q}(z-t_{q})}\mathrm{e}^{\mathrm{i}r_{q+1}(z-t_{q+1})}\cdots\mathrm{e}^{\mathrm{i}r_{2q}(z-t_{2q})}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\sum_{i=1}^{d}\alpha_{i}g_{i}\ast F_{n}^{2q,P}(z),$
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 8. Since $P$ is a convex polyhedron, the last equality is by Lemma 9. For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable [17, Lemma 10]. $\Box$
[…]
Proof  Let $r^{\prime}\in Q_{j}^{m}$. For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=m+r_{j+1}^{\prime}+\cdots+r_{i}^{\prime}$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$. Therefore, we have
$\bigcup_{j=0}^{2q}Q_{j}^{m}$
$=\bigcup_{j=0}^{2q}\bigg{\{}r\in\mathbb{R}^{2q}\,\bigg{|}\,r_{i}=\sum_{l=j+1}^{i}r_{l}^{\prime}+m\ (j<i\leq 2q),\ r_{i}=-\sum_{l=i+1}^{j}r_{l}^{\prime}+m\ (0\leq i<j),$
$\qquad\qquad 0\leq r_{i}\leq m\ (i=0,\ldots,j-1,j+1,\ldots 2q)\}$
[…]
=$=\bigg{\{}r\in\mathbb{R}^{2q}\,\bigg{|}\,\bigg{|}\sum_{i=l}^{k}r_{i}\bigg{|}\leq 1,\ l\leq k\bigg{\}},$
which completes the proof of the lemma. $\Box$
Proposition 11    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{poly},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=1}^{n}\big{(}(R_{n}(x_{i,j})^{*})^{q}\big{)}_{j^{\prime},m}\big{(}R_{n}(x_{i,l})^{q}\big{)}_{m,l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}d_{l}(z)$
$=\sum_{i=1}^{d}\alpha_{i}\sum_{m=1}^{n}\bigg{|}\sum_{j=1}^{N}\sum_{j^{\prime}=1}^{n}d_{j}(z)\big{(}(R_{n}(x_{i,j}))^{q}\big{)}_{m,j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}\geq 0$
for $z\in\mathbb{T}$. In addition, let $\tilde{x}_{n,j}=S_{n}(R_{n}(x_{j}))$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{sep},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{2q-1}=1}^{n}$
$\big{(}(R_{n}(a)^{*})\big{)}_{j^{\prime},m_{1}}\big{(}(R_{n}(a)^{*})\big{)}_{m_{1},m_{2}}\cdots\big{(}(R_{n}(a)^{*})\big{)}_{m_{q-1},m_{q}}\big{(}R_{n}(a)\big{)}_{m_{q},m_{q+1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{2q-1},l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)$
$\qquad\sum_{m_{q}=1}^{n}\bigg{|}\sum_{j^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{q-1}=1}^{n}\big{(}R_{n}(a)\big{)}_{m_{q},m_{1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{q-2},m_{q-1}}\big{(}R_{n}(a)\big{)}_{m_{q-1},j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}$
$\geq 0.$
Thus, $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite. $\Box$
Proposition 12    Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
$\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)\mathrm{d}t$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
$\geq 0$
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},x_{2}],[y_{1},y_{2}])\mapsto\overline{\tilde{k}_{1,j}(x_{1},y_{1})}\tilde{k}_{2,j}(x_{2},y_{2})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$. $\Box$
Lemma 13    We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
Proof  The number of terms in $F_{n}^{q,P}$ is equal to the number of terms in $\sum_{m,l=0}^{n-1}(T_{1}\cdots T_{2q})_{m,l}$, where $T_{1},\ldots,T_{2q}$ are Toeplitz matrices. Thus, we have
$|F_{n}^{q,P}(z)|\leq\frac{1}{n}n^{q+1}=n^{q}.$
$\Box$
Proposition 16    Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
Proof  We first show $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)=\sup_{a\in\mathcal{S}}a(z)$ for $\mathcal{S}\subseteq\mathcal{A}_{+}$ for any $z\in\mathbb{T}$. Let $a\in\mathcal{S}$. Since $a(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$, we have $a\leq_{\mathcal{A}}b$, where $b\in\mathcal{A}$ is defined as $b(z)=\sup_{a\in\mathcal{S}}a(z)$. Thus, $b$ is an upper bound of $\mathcal{S}$, and we have $\sup^{\mathcal{A}}_{a\in\mathcal{S}}a\leq_{\mathcal{A}}b$, which means $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$. Conversely, since $a\leq_{\mathcal{A}}\sup^{\mathcal{A}}_{a\in\mathcal{S}}a$ for $a\in\mathcal{S}$, we have $a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$ for any $z\in\mathbb{T}$. Thus, we have $\sup_{a\in\mathcal{S}}a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$.
Therefore, we have
$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$ $=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}}\!^{\mathcal{A}}\bigg{|}\sum_{i=1}^{N}f(x_{i})^{*}\sigma_{i}\bigg{|}_{\mathcal{A}}\bigg{]}(z)=\frac{1}{N}\mathrm{E}\bigg{[}\bigg{(}\sup_{f\in\mathcal{F}}\!^{\mathcal{A}}\bigg{|}\sum_{i=1}^{N}f(x_{i})^{*}\sigma_{i}\bigg{|}_{\mathcal{A}}\bigg{)}(z)\bigg{]}$
$=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}\overline{f(x_{i})}\sigma_{i}\bigg{|}\bigg{]}=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}f(x_{i})\sigma_{i}\bigg{|}\bigg{]}=\hat{R}_{N}(\mathbf{x},\mathcal{F}(z)),$
where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued. $\Box$
Theorem 15    Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep}q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$
for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
According to Hashimoto et al. [14], the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
Lemma 17
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq C\}$. Then, we have
$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})\leq_{\mathcal{A}}\frac{B}{N}\bigg{(}\sum_{i=1}^{N}\|k(x_{i},x_{i})\|_{\mathcal{A}}\bigg{)}^{1/2}1_{\mathcal{A}}.$
We obtain the following lemma directly by Lemma 4.2 in [31].
Lemma 18
Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. Assume also for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have
$\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(x_{i})(z),y_{i}(z))\sigma_{i}\bigg{]}=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\sum_{i=1}^{N}g(f(x_{i}),y_{i}(z))\sigma_{i}\bigg{]}\leq L\hat{R}_{N}(\mathbf{x},\mathcal{F}(z)).$
We apply the following lemma to obtain the inequality (3).
Lemma 19
Let $g$ be the same map defined in Lemma 18. For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)](z)\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
Proof  With probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)](z)=\mathrm{E}[g(f(X)(z),Y(z))]$
$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}$
$\quad=\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
The inequality follows by Theorem 3.1 in [31]. $\Box$
The following lemma completes the proof of Theorem 15.
Lemma 20
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3. Then, for $x\in\mathcal{A}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
Proof  We evaluate $\|k_{n}(x,x)\|_{\mathcal{A}}$. For $k_{n}^{\operatorname{poly},q}$, we have
$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{\|}_{\operatorname{op}}$
$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x)\|^{2q}_{\operatorname{op}}.$
The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
$|S_{n}(A)(z)|$ $=\bigg{|}\frac{1}{n}\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}\bigg{|}$
$=\bigg{|}\operatorname{tr}\bigg{(}\frac{1}{n}\mathbf{1}\operatorname{diag}(\mathrm{e}^{0},\ldots,\mathrm{e}^{\mathrm{i}(n-1)})A\operatorname{diag}(\mathrm{e}^{0z},\ldots,\mathrm{e}^{\mathrm{i}(n-1)z})\bigg{)}\bigg{|}$
$\leq\frac{1}{n}\|\mathbf{1}\|_{1}\|\operatorname{diag}(\mathrm{e}^{0},\ldots,\mathrm{e}^{\mathrm{i}(n-1)})A\operatorname{diag}(\mathrm{e}^{0z},\ldots,\mathrm{e}^{\mathrm{i}(n-1)z})\|_{\operatorname{op}}\leq\|A\|_{\operatorname{op}}$
for any $z\in\mathbb{T}$. Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$. We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ [32, Chapter IX, Section 2].
In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$. Thus, we have
$\|R_{n}(x)\|_{\operatorname{op}}$ $=\|R_{n}(x)v\|=\|Q_{n}^{*}M_{x}Q_{n}v\|=\bigg{\|}Q_{n}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}\leq\bigg{\|}Q_{n+1}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}$
$\leq\|Q_{n+1}^{*}M_{x}Q_{n+1}\|_{\operatorname{op}}=\|R_{n+1}(x)\|_{\operatorname{op}},$
where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function. $\Box$
Appendix B Details of Remark 10
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 9 from the perspective of $q$. The following lemma is by Brandlini and Travaglini [25].
Lemma 21
Let $H_{n}^{q,P}(t)=n^{q}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{d}r\mathrm{d}j$ for $t\in\mathbb{T}^{q}$, where $\chi_{S}$ is the characteristic function with respect to a set $S$. Then, we have $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$.
Proof  Let $u_{j,t}(r)=\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}$. Then, we have
$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j,$
where $\hat{\cdot}$ is the Fourier transform, and the third equality is by the Poisson summation formula. $\Box$
To understand the convrgence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 9, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 [25, Theorem 1]. Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
Lemma 22
For $z\in\mathbb{T}$, we have
$\lim_{n\to\infty}g\ast H_{n}^{q,P}(z)-g(z)=\int_{0}^{1}\lim_{n\to\infty}\bigg{(}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\bigg{)}\mathrm{d}j.$
Moreover, let $h^{P}_{s}(z)=\int_{t\in\mathbb{T}^{q}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$. Let $q=2$, $Q=\{(r_{1},-r_{2})\,\mid\,(r_{1},r_{2})\in P\}$, $P_{1}^{\prime}=[-1,1]$, and $P_{2}=\{(r_{1},r_{2})\,\mid\,|r_{1}|\leq 1,\ |r_{2}|\leq 1,\ |r_{1}+r_{2}|\leq 1,\ |r_{1}-r_{2}|\leq 1\}$. Let $g(z)=\overline{x_{1}(z_{1})}\overline{x_{2}(z_{2})}y_{1}(z_{3})y_{4}(z_{4})$ for $z=[z_{1},z_{2},z_{3},z_{4}]$. Then, for $z\in\mathbb{T}^{2}$, there exists $C(z)\geq 0$ such that we have
$\frac{1}{2}|h^{P}_{jn}(z)+h^{Q}_{jn}(z)|\leq\frac{1}{2}C(z)\sum_{j=1}^{2}$ $\bigg{|}\int_{t\in\mathbb{T}}\int_{jnP_{1}^{\prime}}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{j}(t_{j})\mathrm{d}r\mathrm{d}t-x_{j}(z_{j})\bigg{|}$
$+\frac{1}{2}\bigg{|}\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t_{1}-g(z)\bigg{|}.$ (6)
Proof  We have
$n^{q}\int_{t\in\mathbb{T}^{q}}\int_{0}^{1}\int_{jP}\mathrm{e}^{\mathrm{i}nr\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}j\mathrm{d}t-g(z)=\int_{0}^{1}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j.$
Let ${h}^{P}_{s}(z)=\int_{t\in\mathbb{T}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$. Then, $h^{P}_{0}(z)=0$, and since $\lim_{s\to\infty}h^{P}_{s}(z)=0$, there exists $D(z)>0$ such that $|h^{P}_{s}(z)|\leq D(z)$ for any $s>0$. Thus, $|h^{P}_{jn}(z)|\leq D(z)$ for any $n\in\mathbb{N}$ and $0<j\leq 1$. Thus, by the bounded convergence theorem, we have
$\lim_{n\to\infty}\int_{0}^{1}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j=\int_{0}^{1}\lim_{n\to\infty}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j.$
In addition, let $P_{1}=[-1,1]^{2}$ and $P_{2}=\{(r_{1},r_{2})\,\mid\,|r_{1}|\leq 1,\ |r_{2}|\leq 1,\ |r_{1}+r_{2}|\leq 1,\ |r_{1}-r_{2}|\leq 1\}$. Then, we have
$h^{P}_{jn}(z)+h^{Q}_{jn}(z)=h^{P_{1}}_{jn}(z)+h^{P_{2}}_{jn}(z).$
For $h^{P_{1}}_{jn}(z)$, we have
$|h^{P_{1}}_{jn}(z)|$ $=\bigg{|}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{2}-t)}x_{2}(t)\mathrm{d}r\mathrm{d}t-g(z)\bigg{|}$
$\leq\bigg{|}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t\bigg{(}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{2}-t)}x_{2}(t)\mathrm{d}r\mathrm{d}t-x_{2}(z_{2})\bigg{)}\bigg{|}$
$\quad+\bigg{|}x_{2}(z_{2})\bigg{(}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t-x_{1}(z_{1})\bigg{)}\bigg{|}.$
In the same manner as $h_{s}^{P}(z)$, there exists $\tilde{C}(z)>0$ such that $|\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t|\leq\tilde{C}(z)$. By setting $C(z)=\max\{\tilde{C}(z),|x_{2}(z_{2})|\}$, we obtain the result. $\Box$
The left hand side is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$. The first term in the right hand side of Eq. (6) is described by the convergence of $h_{jn}^{[-1,1]}(z)$, which corresponds to the one-dimensional ($(q-1)$-dimensional) Fejér kernel. Since we have an additional term $1/2|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$. Based on this observation, we expect that the convergence in Theorem 7 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.
Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. [33, Proposition 4.5] showed an approximate representer theorem for RKHMs over von Neumann-algebras. Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
Proposition 23
Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$. Let $g:\mathcal{A}^{N}\to\mathcal{A}_{+}$ be a continuous map, let $h:\mathcal{A}_{+}\to\mathcal{A}_{+}$ satisfy $h(c)\leq_{\mathcal{A}}h(d)$ for $c,d\in\mathcal{A}_{+}$ with $c\leq_{\mathcal{A}}d$. Let $L(f)=g(f(x_{1}),\ldots,f(x_{N}))+h(|f|_{k})$ for $f\in\mathcal{M}_{k}$. If there exists a solution $f_{0}$ of the minimization problem $\min_{f\in\mathcal{M}_{k}}L(f)$, then for any $\epsilon>0$, there exists $\tilde{f}$ that admits the representation of $\tilde{f}=\sum_{i=1}^{N}\phi(x_{i})c_{i}$ and $\|L(\tilde{f})-L(f_{0})\|\leq\epsilon$.
Proof  For $f_{1},f_{2}\in\mathcal{M}_{k}$, let $\theta_{f_{1},f_{2}}:\mathcal{M}_{k}\to\mathcal{M}_{k}$ defined as $\theta_{f_{1},f_{2}}(v)=f_{1}\left\langle f_{2},v\right\rangle_{k}$ for $v\in\mathcal{M}_{k}$. Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, and let $\mathcal{B}=\overline{\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$. In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see [12, p8]). Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), it has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ [23, Theorem I.4.8]. For each $v\in\mathcal{N}_{k}$, we have
$\|v-b_{i}v\|_{k}^{2}$ $=\|\left\langle(1-b_{i})v,(1-b_{i})v\right\rangle_{k}\|_{\mathcal{A}}=\|\theta_{(1-b_{i})v,(1-b_{i})v}\|_{\mathcal{L}(\mathcal{M}_{k})}$
$=\|(1-b_{i})\theta_{v,v}(1-b_{i})^{*}\|_{\mathcal{L}(\mathcal{M}_{k})}=\|(1-b_{i})\theta_{v,v}(1-b_{i})\|_{\mathcal{L}(\mathcal{M}_{k})}.$
Thus, we have $\lim_{i}b_{i}v=v$. Since $g$ is continuous, we have
$\lim_{i}g(\left\langle\phi_{k}(x_{1}),b_{i}f_{0}\right\rangle_{k},\ldots,\left\langle\phi_{k}(x_{N}),b_{i}f_{0}\right\rangle_{k})=\lim_{i}g(\left\langle b_{i}^{*}\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle b_{i}^{*}\phi_{k}(x_{N}),f_{0}\right\rangle_{k})$
$=\lim_{i}g(\left\langle b_{i}\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle b_{i}\phi_{k}(x_{N}),f_{0}\right\rangle_{k})=g(\left\langle\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle\phi_{k}(x_{N}),f_{0}\right\rangle_{k}).$
Therefore, for $\epsilon>0$, there exists $i$ such that
$\|g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N}))\|_{\mathcal{A}}\leq\epsilon.$
Since $b_{i}\geq_{\mathcal{L}(\mathcal{M}_{k})}0$, there exists $c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$ such that $b_{i}=c^{2}$. Thus, $b_{i}-b_{i}^{2}=c^{2}-c^{4}=c(1-c^{2})c=c^{*}(1-b_{i})c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$. Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}$. As a result, we obtain
00 0 $\leq_{\mathcal{A}}L(b_{i}f_{0})-L(f_{0})$
$=g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))+h(|b_{i}f_{0}|_{k})-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N}))-h(|f_{0}|_{k})$
$\leq_{\mathcal{A}}g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N})).$
Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition. $\Box$
Appendix D Experimental Details and Additional Results
We provide experimental details below. All the experiments in this paper were executed with Python 3.9 on an Intel(R) Core(TM) i9-10885H 2.4GHz processor with the Windows 10 operating system.
Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order. (Average value of results of five different runs. The error bar represents the standard deviation.)
D.1 Experiment with synthetic data
We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$. We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. For $x^{i}$, $y^{i}$, $\tilde{x}^{i}$, and $a$, we discretized each function with 30 equally spaced points on $\mathbb{T}$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$ discussed in Remark 14, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 3 shows the result. The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
D.2 Experiment with MNIST
We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.

### v3

@@
 Abstract
-In this paper, we propose a new class of positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra.
-We focus on kernels whose inputs and outputs are functions and generalize existing kernels, such as polynomial, product, and separable kernels, by introducing a truncation parameter $n$ that describes the noncommutativity of the products appearing in the kernels.
-When $n$ goes to infinity, the proposed kernels tend to the existing commutative kernels.
-If $n$ is finite, they exhibit different behavior, and the noncommutativity induces interactions along the data function domain.
-We show that the truncation parameter $n$ is a governing factor leading to performance enhancement: by setting an appropriate $n$, we can balance the representation power and the complexity of the representation space.
-The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels.
+$C^{*}$-algebra-valued kernels could pave the way for the next generation of kernel machines.
+To further our fundamental understanding of learning with $C^{*}$-algebraic kernels, we propose a new class of positive definite kernels based on the spectral truncation.
+We focus on kernels whose inputs and outputs are vectors or functions and generalize typical kernels by introducing the noncommutativity of the products appearing in the kernels.
+The noncommutativity induces interactions along the data function domain.
+We show that it is a governing factor leading to performance enhancement: we can balance the representation power and the model complexity.
+We also propose a deep learning perspective to increase the representation capacity of spectral truncation kernels.
+The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
 1 Introduction
-Kernel methods have been one of the most fundamental tools in machine learning [1, 2, 3, 4, 5].
+Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf & Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017).
 They have been applied, for example, to ridge regression, principal component analysis, and support vector machine.
-Kernel methods are characterized by function spaces called reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels.
+Kernel methods are characterized by reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels.
 Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel.
-Product kernels, which are constructed by the product of multiple kernels, have also been considered [1, 6].
-Standard positive definite kernels are real- or complex-valued, and are well-suited to learn scalar-valued functions.
-Kernel methods for vector-, function-, or operator-valued outputs have also been investigated [4, 7, 8].
-The kernel, in this case, is instead an operator- or function-valued positive definite function, and the associated feature space is a vector-valued RKHS (vvRKHS) [9, 7, 10] or a reproducing kernel Hilbert $C^{*}$-module (RKHM) [8].
-RKHM is a generalization of RKHS by means of $C^{*}$-algebra, and it also generalizes vvRKHS.
-$C^{*}$-algebra is a generalization of the space of complex values and has a structure of the norm, product, and involution [11, 12].
+Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf & Smola, 2001; Thomas, 2008).
+Standard positive definite kernels are scalar-valued, and are well-suited to learn scalar-valued functions.
+Kernel methods for vector- and function-valued outputs have also been investigated (Álvarez et al., 2012; Kadri et al., 2016).
+The kernels, in these cases, are instead operator-valued, and the associated feature space is vector-valued RKHSs (vvRKHSs) (Kadri et al., 2012, 2016; Minh et al., 2016).
+There are at least two challenges for vvRKHS methods: the computational cost and choice of kernels.
+A typical kernel is the separable kernel, which is defined by the product of a scalar-valued kernel and a positive semi-definite operator (Álvarez et al., 2012).
+Although applying the separable kernel is computationally efficient, it identifies dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables.
+Several attempts have been made to construct kernels that go beyond separable kernels.
+Huusari & Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement.
+Hashimoto et al.
+(2023a) proposed to use the product of circulant matrices and general squared matrices to construct kernels.
+Using this kernel, one can generalize the convolution and capture the effect of interactions of different Fourier components.
+However, if we need an $m$-dimensional vector-valued outputs with nonseparable kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general.
+Thus, with vvRKHSs, we cannot resolve the two issues at the same time.
+In this work, we address these two challenges by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021).
+RKHM is a generalization of RKHS by means of $C^{*}$-algebra.
+$C^{*}$-algebra is a generalization of the space of complex values and has structures of the norm, product, and involution (Murphy, 1990; Lance, 1995).
 It unifies operators and functions.
-Thus, the operator- and function-valued kernels for vvRKHSs and RKHMs can be unified as $C^{*}$-algebra-valued kernels.
-It has been shown that the norms in $C^{*}$-algebras alleviate the dependency of the generalization bound on the output dimension [13] and the product structure of a $C^{*}$-algebra goes beyond the convolution operation [14].
-One big challenge for kernel methods is the choice of positive definite kernels.
-Especially for vvRKHSs and RKHMs, the construction of good $C^{*}$-algebra-valued kernels is not obvious.
-A typical kernel is the separable kernel, which is defined by the product of a real- or complex-valued kernel and a positive semi-definite operator [4].
-Attempts have been made to construct positive definite kernels that go beyond separable kernels.
-Huusari and Kadri [15] proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement.
-Hashimoto et al.
-[14] proposed to use the product of circulant matrices and general squared matrices to construct positive definite kernels.
-By using the kernel, we can generalize the convolution and capture the effect of interactions of different Fourier components.
-In this paper, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the field of noncommutative geometry and $C^{*}$-algebra [16, 17, 18].
-We focus on the case of functional inputs and outputs, and consider the $C^{*}$-algebra of continuous functions.
-We generalize typical existing kernels: polynomial, product, and separable kernels using the pair of maps called the approximate order isomorphism, which takes an important role in the spectral truncation [19, 17, 20].
-The proposed kernels are parameterized by a natural number $n$ that corresponds to the dimension of the truncated space.
-We approximate functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices.
-Thus, $n$ describes the resolution of the discretization of the function, and we call it the truncation parameter.
-We show that the proposed kernels converge to the existing kernels as $n$ goes to infinity.
-To analyze functional data, we have been focused on how to interpolate observed data to obtain functions [21, 22, 8].
-On the other hand, our results shed light on good effects of the discretization (setting $n$ as a finite number) on learning process.
-Indded, the truncation parameter $n$ takes an important role from at least two perspectives.
-First, one of the important structures in the existing kernels is the product, and $n$ describes the noncommutativity of the product.
-Although the product of two functions is too simple to capture structures of these functions, the product of two approximated Toeplitz matrices is more complex, which enables us to capture the structures of the functions.
-More precisely, for two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$.
+In this framework, kernels are generalized to $C^{*}$-algebra-valued kernel functions and allow us to consider function-valued kernels, leading to function-valued Gram matrices.
+By evaluating the values of the function-valued Gram matrix at $m$ different points, we obtain $m$ scalar-valued Gram matrices.
+This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels.
+To obtain kernels going beyond the separable kernels, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes & van Suijlekom, 2021).
+Our class of kernels is large and encompasses previous commutative kernel functions such as polynomial, product, and separable kernels.
+The proposed kernels are parameterized by a natural number $n$ corresponding to the dimension of the truncated space.
+They can be applied to both vector and functional inputs.
+For vector inputs, we regard the elements of them as the values of functions.
+We approximate the input functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices, whose $(i,j)$-entry depends only on $i-j$.
+Thus, $n$ describes the resolution of the discretization, and we call it the truncation parameter.
+We show that the proposed kernels converge to the commutative kernels as $n$ goes to infinity.
+In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process.
+The truncation parameter $n$ plays an important role from at least two perspectives.
+First, an important structure in typical commutative kernels is the product, and $n$ describes the noncommutativity of the product.
+For two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$.
 Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product.
+This situation corresponds to the case of $n=\infty$.
 On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$.
 This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$.
-Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHM.
+This corresponds to the case $n<\infty$.
+Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity.
 We show that if $n$ is small, then the representation power is low, and the model complexity is small.
 On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large.
-The expectation loss for learning a function is decomposed into the empirical loss, controlled by the representation power, and the complexity.
-By introducing the parameter $n$ and setting $n$ to balance the empirical loss and the complexity, we obtain higher performance compared to the case of existing kernels.
+The expectation loss for learning a function is controlled by the representation power and the complexity.
+By introducing the parameter $n$ and setting $n$ to balance them, we obtain higher performance compared to the typical commutative kernels.
+In addition, to resolve the issue of the interactions becoming weaker as $n$ becomes larger (i.e., representation power becomes higher), we propose a deep approach to obtain models with higher representation powers while keeping the interactions.
 Our contributions are summarized as follows:
-  • We propose a new class of function-valued positive definite kernels that generalize existing kernels: polynomial, product, and separable kernels.
-The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter $n$, which allows us to introduce noncommutativity into the learning process.
-The proposed kernels induce the interactions along the data function domain.
-(Definition 6 and Subsection 3.1)
+  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that generalize typical kernels.
+The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which induces the interactions along the data function domain by introducing the noncommutativity into the learning process.
+(see Definition 6 and Subsection 3.1)
   • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels.
-The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHMs.
-(Section 4)
-  • We apply the proposed kernels to kernel ridge regression. The result illustrates the advantage of introducing the truncation parameter $n$ in practical settings. (Section 5)
-We emphasize that although we focus on the case where the inputs and outputs are functions and three types of kernels, our idea of constructing new positive definite kernels is general.
-We establish a new framework that allows us to go beyond the usual commutative kernels and to assign a significance to noncommutative structures by discretizing functions.
+The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity.
+(see Section 4)
+  • We propose a deep approach to gain the representation power while keeping the noncommutativity in the kernels. (see Section 6)
+Our framework allows us to go beyond the usual commutative kernels and to develop a new perspective on the significance of noncommutative structures in learning systems.
 2 Preliminaries
-We review mathematical notions required in this paper.
 2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
@@
   2. 2.
-    $\mathcal{A}$ is a normed space endowed with $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds.
+    $\mathcal{A}$ is a normed space endowed with a norm $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds.
 In addition, $\mathcal{A}$ is complete with respect to $\|\cdot\|_{\mathcal{A}}$.
@@
     For $a\in\mathcal{A}$, the $C^{*}$ identity $\|a^{*}a\|_{\mathcal{A}}=\|a\|_{\mathcal{A}}^{2}$ holds.
-If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit, and denoted by $1_{\mathcal{A}}$.
+If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit and denoted by $1_{\mathcal{A}}$.
 In this paper, we focus on the $C^{*}$-algebra of continuous functions.
@@
 Then, $\mathcal{A}:=C(\mathbb{T})$ is a $C^{*}$-algebra by means of the product: $(cd)(z)=c(z)d(z)$, involution: $c^{*}(z)=\overline{c(z)}$, and norm: $\|c\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}}|c(z)|$ for $c,d\in\mathcal{A}$.
-We review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
+The unit is the constant function $1_{\mathcal{A}}\equiv 1$.
+We now review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
 Definition 2 (Positive)
-An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds.
-For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive.
-We denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
+An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds. For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive.
 Definition 3 (Infimum and minimum)
@@
 We can define an $\mathcal{A}$-valued map $\big{\langle}\cdot,\cdot\big{\rangle}_{k}:\mathcal{M}_{k,0}\times\mathcal{M}_{k,0}\to\mathcal{A}$ as
-$\bigg{\langle}\sum_{i=1}^{n}\phi(x_{i})c_{i},\sum_{j=1}^{l}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{n}\sum_{j=1}^{l}c_{i}^{*}k(x_{i},y_{j})d_{j},$
+$\bigg{\langle}\sum_{i=1}^{N}\phi(x_{i})c_{i},\sum_{j=1}^{M}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i}^{*}k(x_{i},y_{j})d_{j},$
 which enjoys the reproducing property $\left\langle\phi(x),f\right\rangle_{k}=f(x)$ for $f\in\mathcal{M}_{k,0}$ and $x\in\mathcal{X}$.
@@
 The reproducing kernel Hilbert $\mathcal{A}$-module (RKHM) $\mathcal{M}_{k}$ associated with $k$ is defined as the completion of $\mathcal{M}_{k,0}$ with respect to $\|\cdot\|_{k}$.
-See, for example, the references [12, 11, 23, 8] for more details about $C^{*}$-algebra and RKHM.
+See, for example, the references (Lance, 1995; Murphy, 1990; Hashimoto et al., 2021) for more details about $C^{*}$-algebra and RKHM.
 2.2 Spectral truncation on the torus
@@
 Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$.
-Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square integrable functions on $\mathbb{T}$.
+Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square-integrable functions on $\mathbb{T}$.
 Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$.
@@
 $R_{n}(x)_{j,l}=\left\langle e_{j},M_{x}e_{l}\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}}x(t)\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{d}t.$
-Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix [24].
+Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix (Gray, 2006).
 It is characterized only by a vector $r_{n}(x)\in\mathbb{C}^{2n-1}$, where the $(j-l)$th element of $r_{n}(x)$ is $R_{n}(x)_{j,l}$.
 Note that $R_{n}(x)_{j,l}$ is regarded as the $(j-l)$th Fourier component of $x$, and the vector $r_{n}(x)$ is the coordinate of the function $x$ projected on the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$.
-For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=1/n\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$.
+For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=(1/n)\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$.
 The map $S_{n}$ takes the representation matrix $R_{n}(x)$ back to a function that approximates the original function $x$.
@@
 $S_{n}(R_{n}(x))(z)=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j,l=0}^{n-1}\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{e}^{\mathrm{i}(j-l)z}\mathrm{d}t=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}l(z-t)}\mathrm{d}t=x\ast F_{n}(z),$
-where $F_{n}(t)=1/n\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution.
-By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ [25, 17].
+where $F_{n}(t)=(1/n)\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution.
+By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ (Brandolini & Travaglini, 1997; van Suijlekom, 2021).
 Proposition 5
 For each $z\in\mathbb{T}$, $x\ast F_{n}(z)\to x(z)$ as $n\to\infty$.
-The pair of maps $R_{n}$ and $S_{n}$ is called the approximate order isomorphism.
+A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini & Travaglini, 1997).
+Let $P$ be a polyhedron and $jP=\{jz\,\mid\,z\in P\}$ for $j\in\mathbb{Z}$.
+For $t\in\mathbb{T}^{q}$, the Fejér kernel on $\mathbb{T}^{q}$ is defined as
+$F_{n}^{q,P}(t)=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}.$ (1)
 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
-In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. The following three kinds of $\mathcal{A}$-valued kernels are often considered in existing studies [14, 13, 1, 9, 4].
+In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$.
+To obtain vector- or functional-valued outputs, applying vvRKHSs has been investigated.
+However, as we stated in Section 1, to go beyond the separable kernel and reach higher performance, the computational cost scales as $O(m^{3}N^{3})$ for $m$-dimensional outputs with $N$ samples.
+To resolve this issue, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs.
+As we will discuss in the last part of Section 5, the computational cost depends on $mN^{3}$ if we use an $\mathcal{A}$-valued kernel.
+Hashimoto et al.
+(2021) proposed to use the $\mathcal{A}$-valued kernel $k$ where $k(\cdot,\cdot)(z)$ is a complex-valued positive definite kernel for all $z\in\mathbb{T}$.
+Combining it with typical existing kernels (Hashimoto et al., 2023a; Schölkopf & Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012), we obtain the following three kinds of $\mathcal{A}$-valued kernels.
 Example 2
@@
 In addition, $a_{j}\in\mathcal{A}$ is the parameter of $k^{\operatorname{sep},q}$.
-We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the existing kernels.
+We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the typical kernels in Example 2.
 The kernels in Example 2 involve the product of functions in $\mathcal{A}$.
@@
 Definition 6
-With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, let
+With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$, let
 $k_{n}^{\operatorname{poly},q}(x,y)$ $=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)},$
@@
 Here, $A^{*}$ for a matrix $A$ is the adjoint, and we denote by $\tilde{k}_{i,j}(x,y)$ the map $z\mapsto\tilde{k}_{i,j}(x(z),y(z))$.
+Remark 7
+Although the inputs of the kernels in Example 2 and Definition 6 are functions, we can also deal with vector inputs.
+For vector inputs, we can regard them as values of functions, and approximate the integral by the discrete sum.
 In Definition 6, to construct the kernels, we first project functions in $\mathcal{A}$ onto the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ using the map $R_{n}$ and obtain Toeplitz matrices.
@@
 By the definitions of $S_{n}$ and $R_{n}$, we can show the following identity for $x_{1},\ldots,x_{q},y_{1},\ldots,y_{q}\in\mathcal{A}$ (see the proof of Theorem 8 in Appendix A):
-$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (1)
+$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\!\!\!\!\!\!\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (2)
 […]
-$F_{n}^{q,P}(t)=1/n\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}$ (2)
-for $t\in\mathbb{T}^{q}$ is the Fejér kernel on $\mathbb{T}^{q}$ with respect to the polyhedron $P$.
-Here, for $j\in\mathbb{Z}$, $jP=\{jz\,\mid\,z\in P\}$.
-The generalization of the Fejér kernel on $\mathbb{T}$, which is defined in Subsection 2.2, to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated [26, 25].
-Note that $F_{n}^{q,P}(t)$ is a real-valued function since for $r\in P$, we have $-r\in P$.
-Thus, if a kernel in Example 2 is real-valued, then the corresponding kernel in Definition 6 is also real-valued.
-Eq.
-(1) implies that the value of $S_{n}\big{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y_{j})\big{)}$ at $z$ is determined not only by $x_{1}(z),\ldots,x_{q}(z),y_{1}(z),\ldots,y_{q}(z)$, but also by $x_{1}(t_{1}),\ldots,x_{q}(t_{q}),y_{1}(t_{q+1}),\ldots,y_{q}(t_{2q})$ with different values $t_{1},\ldots,t_{2q}$ from $z$.
 3.1 Convergence and interactions
-We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2, respectively.
-In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of existing kernels.
+We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2 as $n$ goes to infinity, respectively.
+In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of typical kernels.
 In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$.
@@
 (2).
-Lemma 9 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini and Travaglini [25].
+Lemma 10 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini & Travaglini (1997).
 We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq.
-(1), and apply Lemmas 8 and 9 to show Theorem 7.
+(2), and apply Lemmas 9 and 10 to show Theorem 8.
 See Appendix A for more details.
@@
 3.2 Positive definiteness
-To define an RKHM introduced in Subsection 2.1, the kernel $k$ has to be positive definite.
+To construct an RKHM, which is introduced in Subsection 2.1, the kernel $k$ has to be positive definite.
 Thus, we investigate the positive definiteness of the proposed kernels.
@@
 Proposition 12
-The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
-As for $k_{n}^{\operatorname{prod},q}(x,y)$, $x$ and $y$ depend on $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products.
+The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
+As for $k_{n}^{\operatorname{prod},q}(x,y)$, it depends on $x$ and $y$ through $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products.
 Thus, we cannot show its positive definiteness.
-However, we can modify it to become positive definite as follows.
+However, we can modify the kernel to become positive definite as follows.
 Proposition 13
-Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
+Let $\beta_{n}\geq-\min_{z\in\mathbb{T}^{q}}F_{n}^{2q,P}(z)$. Then, $\hat{k}_{n}^{\operatorname{prod},q}$ defined below is positive definite.
 $\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
-Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
 To set the value of the parameter $\beta_{n}$, we have the following bound.
@@
 We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness.
-However, the bound in Lemma 13 may not be tight, and even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite.
-In addition, considering non-positive kernels has also been investigated [27, 28].
-Indeed, in practical computations in Subsection 5.1, the algorithms can work even if we set $\beta_{n}<n^{q}$.
+However, even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite.
+In addition, considering non-positive kernels has also been investigated (Ong et al., 2004; Canu et al., 2005).
+Indeed, in practical computations in Subsection 7, the algorithms can work even if we set $\beta_{n}<n^{2q}$.
 Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@
 We now focus on what impact the truncation parameter $n$ has on generalization.
-We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels.
-To derive the generalization bound, we derive an upper bound of the Rademacher complexity.
-Let $\Omega$ be a probability space with the probability measure $\mu$.
-Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be input and output samples from a distributions of $\mathcal{A}_{0}^{d}$-valued and $\mathcal{A}_{1}$-valued random variables $X$ and $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$).
+We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity associated with the proposed kernels.
+According to Lemma 4.2 in Mohri et al.
+(2012), the generalization bound is described by the Rademacher complexity.
+We apply this lemma and derive a generalization bound.
+Let $\Omega$ be a probability space with a probability measure $\mu$.
+Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$).
 Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$.
-Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d.
-Rademacher variables, which take their values on $\{-1,1\}$.
-For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$.
-Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$.
-An $\mathcal{A}$-valued version of the Rademacher complexity is also defined by Hashimoto et al.
-[14] as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$.
-Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
+Let $\mathrm{E}$ be the Bochner integral on $\Omega$ with respect to $\mu$.
 Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$.
@@
 Theorem 16
-Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ are real-valued.
-Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
+Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are real-valued.
+Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
 Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$.
@@
 for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
-The first term of the bound (3) is the empirical loss, and the second term of the bound represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$ associated with $k_{n}$.
+The first term of the right-hand side of (3) is the empirical loss, and the second term represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$.
 There is a tradeoff between these terms.
-If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$ in $\mathcal{A}$ is larger, $\mathcal{M}_{k_{n}}$ can represent more functions.
+If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ in $\mathcal{A}$ is larger, the outputs of ${k_{n}}$ can represent more functions, which enables us to describe the dependency of outputs on inputs more.
+Indeed, if $n=1$, then the output of $k_{n}$ is always a constant function, where the situation is the same as the existing separable kernel discussed in Section 1.
 Thus, the empirical loss (the first term) can become small if $n$ is large.
 On the other hand, according to Theorem 16, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger.
-An advantage of introducing the parameter $n$ and constructing the new kernels based on the existing kernels in Example 2 is that we can control the empirical loss and the complexity through $n$.
-By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$ small.
-To show Theorem 15, we use the $\mathcal{A}$-valued Rademacher complexity.
-Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity.
-Proposition 16
-Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class.
-Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$.
-Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
+An advantage of introducing the parameter $n$ to construct the new kernels based on the typical kernels in Example 2 is that we can control the empirical loss and the complexity through $n$.
+By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$, the left-hand side of Eq.
+(3), small.
 5 Application to Kernel Ridge Regression
@@
 We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$.
-Consider the typical minimization problem for regression, which is also considered in [14]:
+Consider the typical minimization problem for regression, which is also considered by Hashimoto et al.
+(2023a):
 $\min_{f\in\mathcal{M}_{k}}\bigg{(}\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{\mathcal{A}}^{2}+\lambda|f|_{k}^{2}\bigg{)},$ (4)
@@
 The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$.
-Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ is $O((qn^{2}+mn^{2})N^{2})$.
+Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ is $O((qn^{2}+mn^{2})N^{2})$.
 As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$.
 Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$.
-On the other hand, that for $k^{\operatorname{poly},q}$, $\hat{k}^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ is $O(qmN^{2}+mN^{3})$.
 Note that the cost is linear with respect to $m$.
 This is by virtue of considering function-valued kernels and RKHMs.
-The situation is different if we use a vvRKHS.
-For vvRKHSs, to obtain $m$ values as an output, we generally apply a $\mathbb{C}^{m\times m}$-valued kernel.
-In this case, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$.
-We can reduce the cost with respect to $N$, e.g., by Nyström method [29].
+The situation is different if we use vvRKHSs.
+For vvRKHSs, to obtain $m$ values as an output, we apply a $\mathbb{C}^{m\times m}$-valued kernel.
+If the kernel is nonseparable, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$.
+We can reduce the cost with respect to $N$, e.g., by Nyström method (Drineas & Mahoney, 2005).
 Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel.
 However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
-5.1 Numerical results
+6 Deep model with the spectral truncation kernels
+As we discussed in Section 4, there is a tradeoff between the representation power and the model complexity.
+In fact, there are two directions of the representation power: one is about the dependency of the outputs on the inputs discussed in Section 4, and the other is about the interactions along the data function domain.
+One shortcoming of the spectral truncation kernel is that to obtain a model with a high representation power in the sense of the dependency of the outputs on the inputs, the truncation parameter $n$ have to be large, which results in making the interactions along the data function domain weak.
+To gain the two types of representation powers at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n}^{1},\ldots,k_{n}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
+$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (6)
+We consider the case where $c_{i}^{j}$ in Eq. (6) is represented as a finite sum of functions in $\mathcal{A}$. In the following case, the representation power of this model grows exponentially.
+Proposition 17
+Let $k^{j}_{n}(x,x_{i})=\sum_{l=-(n-1)}^{n-1}\tilde{d}_{i,l}^{j}\mathrm{e}^{\mathrm{i}lz}$ with $\tilde{d}_{i,l}^{j}\in\mathbb{C}$.
+Assume $c_{i}^{j}$ is parameterized as $c_{i}^{j}(z)=\sum_{l=-(n-1)}^{n-1}d_{i,l}^{j}e_{\tau_{j},l}$ with $d_{i,l}^{j}\in\mathbb{C}$ and $\tau_{j}\in\mathbb{R}$.
+Here, $e_{\tau_{j},l}\in C(\mathbb{T})$ satisfies $e_{\tau_{j},l}(z)=\mathrm{e}^{\mathrm{i}\tau_{j}lz}$ for $z\in S$ for a subset $S$ of $\mathbb{T}$.
+Then, for $z\in S$, $f$ in Eq.
+(6) is represented as
+$\sum_{l_{1},\ldots,l_{L}=-(n-1)}^{n-1}\sum_{m_{1},\ldots,m_{L}=-(n-1)}^{n-1}\sum_{i=1}^{N}\tilde{d}_{i,l_{1}}^{1}\cdots\tilde{d}_{i,l_{L}}^{L}{d}_{i,l_{1}}^{1}\cdots{d}_{i,l_{L}}^{L}\mathrm{e}^{\mathrm{i}(l_{1}+\cdots+l_{L}+\tau_{1}m_{1}+\cdots\tau_{L}m_{L})z}.$
+Let $V_{n,L}=\operatorname{Span}\{z\mapsto\mathrm{e}^{\mathrm{i}(l_{1}+\cdots+l_{L})z}e_{\tau_{1},m_{1}}(z)\cdots e_{\tau_{L},m_{L}}(z)\,\mid\,l_{1},\ldots,l_{L},m_{1},\ldots,m_{L}\in\{-(n-1),\ldots,n-1\}\}$.
+If $\{\tau_{1},\ldots,\tau_{L}\}$ is linearly independent over $\mathbb{Z}$, then $V_{n,L}$ is described by $2n+1+(2n+1)^{L}$ oscillating functions.
+In this sense, the representation power of the model $f$ grows exponentially with respect to the number of layers $L$.
+Remark 18
+The exponential growth of the representation power is also observed for neural networks (Hanin & Rolnick, 2019).
+Deep learning with kernels has also been proposed (Laforgue et al., 2019; Hasimoto et al., 2023).
+Whereas the growth is obtained by the composition for these existing frameworks, it is obtained by the product for our case.
+Unlike the case of the existing frameworks, the representation power regarding $\mathcal{A}$ does not become high by the composition of functions.
+Remark 19
+The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k^{j}$.
+Since the Rademacher complexity bound depends on the norm of the kernel, it can also grow exponentially with respect to $L$.
+Thus, for the deep setting, in the same manner as Theorem 16, we still have the tradeoff between the representation power and the model complexity.
+In this case, it is controlled by both $n$ and $L$.
+The exponential growth of the Rademacher complexity bound (thus, the generalization bound) with respect to the number of layers is also observed for neural networks (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018).
+Investigating how we can take advantage of the deep setting in the sense of the generalization is future work.
+7 Numerical results
 We show numerical results that illustrate the effects of the noncommutativity on the learning process.
-5.1.1 Experiment with synthetic data
-To observe the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compare them to the existing kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$), we consider a regression task with synthetic data.
+7.1 Experiment with synthetic data
+We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compared them to the commutative kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$).
+We considered a regression task with synthetic data.
 For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$.
 We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$.
-Here, $\xi_{z,i}$ and $\eta_{z,i}$ are sampled independently from the Gaussian distribution with mean 0 and standard deviation 1.
+Here, $\xi_{z,i}$ and $\eta_{z,i}$ were sampled independently from the Gaussian distribution with mean 0 and standard deviation 1.
 We set the sample size $N$ as 1000.
@@
 Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case.
-For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-|x-y|^{2}}$.
+For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-|x-y|^{2}}$.
 We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$.
-For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1/\Delta^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=\mathrm{e}^{\sin z}$.
+For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1\cdot 2\pi^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=2\pi\mathrm{e}^{\sin z}$.
 We estimated $f$ using kernel ridge regression.
+We applied the cross-validation grid search to find an optimal regularization parameter $\lambda$.
 Figure 2 (a-c) illustrates the test error.
@@
 3 are balanced, and the best possible expected error is obtained.
-We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the existing kernel ($n=\infty$).
+We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the typical commutative kernel ($n=\infty$).
 We have similar results for the other two kernels.
@@
 Thus, the test error becomes large when $n=\infty$.
-In addition, as discussed in Remark 14, $\hat{k}^{\operatorname{prod},q}$ may not be positive definite.
+In addition, as discussed in Remark 15, $\hat{k}_{n}^{\operatorname{prod},q}$ may not be positive definite.
 However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case.
-See Appendix D.1 for more details.
+See Appendix E.1 for more details.
+We also compared the performance and the computational time of RKHMs associated with the proposed kernels to a vvRKHS with the same setting as above.
+For the vvRKHS, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$.
+This nonseparable kernel is proposed by Lim et al.
+(2015).
+Table 1 shows the result.
+We can see that the RKHMs outperform the vvRKHS, and the computational time for the RKHMs are smaller than that for the vvRKHS.
+Note that we have already confirmed the proposed kernels outperform the existing separable kernel $k_{\infty}^{\operatorname{sep},q}$ in the above experiment (See Figure 2 (c)).
+Finally, we observed the behavior of the deep model.
+We generated data in the same manner as above, but we set $N=300$.
+We used the kernel ${k}_{n}^{\operatorname{prod},q}$ and set $L=1,2,3$, $\tau_{1}=\sqrt{2}$, $\tau_{2}=\sqrt{3}$, and $\tau_{3}=\sqrt{5}/2$.
+We set $n$ so that the numbers of parameters are the same for all values of $L$.
+We used the loss function $\|\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{k}\|_{L^{2}(\mathbb{T})}+0.1\||f|_{k}\|_{L^{2}(\mathbb{T})}$ for the deep model $f$ defined in Eq.
+(6).
+Figure 2 (d) shows the result.
+We can see that as the number of layers increases, we can obtain a higher performance model even with the same number of parameters.
 (a) $k={k}_{n}^{\operatorname{poly},q}$
@@
 (c) $k=k_{n}^{\operatorname{sep},q}$
-(d) $k=\hat{k}_{n}^{\operatorname{prod},q}$
+(d) Deep approach
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 2: (a-c) Test error of the regression task with different values of $n$.
-(Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the image recovering task with different values of $n$.
-(Box plot of results of five independent runs with different training and test data.) (e) Output images of the image recovering task with different values of $n$.
-5.1.2 Experiment with MNIST
-We consider an image recovering task with MNIST [30].
+(Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the regression task with deep approach with different $n$ and $L$.
+The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases.
+(e) Output images of the image recovering task with different values of $n$.
+Table 1: Test error and CPU time of the regression task with RKHM and vvRKHS
+Test Error CPU Time (s)
+RKHM, $k=k_{n}^{\operatorname{prod},q}$ ($n=9$) 0.0113$\pm$1.63×$\times 10^{-5}$ 149.3$\pm$2.392
+RKHM, $k=k_{n}^{\operatorname{sep},q}$ ($n=10$) 0.0385$\pm$5.04388×$\times 10^{-5}$ 22.79$\pm$ 0.5742
+vvRKHS 0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
+7.2 Experiment with MNIST
+We consider an image recovering task with MNIST (LeCun et al., 1998).
 For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00.
@@
 We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels.
-We try to estimate a function that transforms an image with the missing part into the original image.
+We tried to estimate a function that transforms an image with the missing part into the original image.
 We applied kernel ridge regression with the regularization parameter $\lambda=0.01$.
-We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-0.1|x-y|^{2}}$.
+We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-0.1|x-y|^{2}}$.
 We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$.
-Figure 2 (d) and (e) show the test error and the output images with different values of $n$.
-We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$.
-As for the output images, when $n=\infty$, i.e., the existing kernel $k^{\operatorname{prod},q}$, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image.
+Figure 2 (e) shows the output images with different values of $n$.
+When $n=\infty$, i.e., the commutative kernel, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image.
 When $n=70$, we can recover the image the clearest.
-6 Generalization to Other $C^{*}$-algebras
-Setting approximate order isomorphisms for more general $C^{*}$-algebras has been investigated.
-Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
-Continuous functions on high-dimensional torus [20]
-[…]
-Continuous functions on the sphere [19]
-For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space.
-Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$.
-We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
 8 Conclusion and Limitation
-In this paper, we proposed a new class of positive definite kernels for functional inputs and outputs based on the spectral truncation.
-The proposed kernels are indexed by a truncation parameter $n$, and they converge to existing kernels: polynomial, product, and separable kernels, as $n$ goes to infinity.
-Setting $n$ as a finite value, we can introduce noncommutativity into the learning process.
-More precisely, we can induce interactions along the data function domain and control the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels.
-Since $n$ describes the resolution of the discretization functions, our results show the positive effect of discretization on generalization.
-We applied the proposed kernel to kernel ridge regression and numerically observed that the proposed kernels outperformed the existing kernels.
+In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation.
+The proposed kernels with the framework of RKHMs resolve two shortcomings of the framework of vvRKHSs for vector- or function-valued outputs at the same time: computational cost and choice of kernels.
+By considering function-valued positive definite kernels, not operator-valued ones, we can alleviate the computational cost.
+At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain.
+We also showed that we can control the tradeoff between the representation power and the model complexity associated with the proposed kernels.
+In addition, we proposed a deep approach to obtain models with higher representation powers.
 In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus.
-Although we can generalize this setting and define positive definite kernels for more general inputs and outputs as we discussed in Section 6, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
+Although this setting includes many important examples, e.g., periodic time-series data, continuous functions defined on a compact set in $\mathbb{R}$, finite-dimensional vectors, investigations for more general $C^{*}$-algebras allow us to apply the proposed kernel more general settings.
+As we discuss in Appendix D, we can generalize this setting and define positive definite kernels for more general inputs and outputs.
+However, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
 Acknowledgements
@@
 If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
-Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)\to k^{\operatorname{poly},q}(x,y)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are the same. We have
-$k_{n}^{\operatorname{poly},q}(x,y)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}$
+Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$.
+The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are similar.
+We have
+$k_{n}^{\operatorname{poly},q}(x,y)(z)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}(z)$
 $=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q}=0}^{n-1}R_{n}(x_{i})^{*}_{j,r_{1}}\cdots R_{n}(x_{i})^{*}_{r_{q-1},r_{q}}R_{n}(y_{i})_{r_{q},r_{q+1}}\cdots R_{n}(y_{i})_{r_{2q-1},l}\mathrm{e}^{\mathrm{i}(j-l)z}$
@@
 $\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
-$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
+$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{2q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
 $\qquad\cdot\mathrm{e}^{\mathrm{i}r_{1}(z-t_{1})}\cdots\mathrm{e}^{\mathrm{i}r_{q}(z-t_{q})}\mathrm{e}^{\mathrm{i}r_{q+1}(z-t_{q+1})}\cdots\mathrm{e}^{\mathrm{i}r_{2q}(z-t_{2q})}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
@@
 In addition, we set $j=r_{0}$ and $j=r_{q}$.
-The second to the last equality is derived by Lemma 8.
-Since $P$ is a convex polyhedron, the last equality is by Lemma 9.
-For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable [17, Lemma 10].
+The second to the last equality is derived by Lemma 9 below.
+Since $P$ is a convex polyhedron, $g_{i}\ast F_{n}^{2q,P}(z)\to g_{i}(z)$ as $n\to\infty$ by Lemma 10.
+For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable (van Suijlekom, 2021, Lemma 10).
 $\Box$
@@
 Proof  Let $r^{\prime}\in Q_{j}^{m}$.
-For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=m+r_{j+1}^{\prime}+\cdots+r_{i}^{\prime}$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$.
+For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=r_{i}^{\prime}+\cdots+r_{j+1}^{\prime}+m$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$.
 Therefore, we have
@@
 which completes the proof of the lemma. $\Box$
-Proposition 11    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
+Proposition 12    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
 Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
@@
 Thus, $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite. $\Box$
-Proposition 12    Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
+Proposition 13    Let $\beta_{n}\geq-\min_{z\in\mathbb{T}^{q}}F_{n}^{2q,P}(z)$. Then, $\hat{k}_{n}^{\operatorname{prod},q}$ defined below is positive definite.
 $\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
-Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
 Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
-$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
-$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
-$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)\mathrm{d}t$
-$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
+$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}\hat{k}_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
+$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
+$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)\mathrm{d}t$
+$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
 $\geq 0$
 for $z\in\mathbb{T}$.
-The last inequality is derived since the map $([x_{1},x_{2}],[y_{1},y_{2}])\mapsto\overline{\tilde{k}_{1,j}(x_{1},y_{1})}\tilde{k}_{2,j}(x_{2},y_{2})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$.
+The last inequality is derived since the map $([x_{1},\ldots,x_{2q}],[y_{1},\ldots,y_{2q}])\mapsto\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{2j-1},y_{2j-1})}\tilde{k}_{2,j}(x_{2j},y_{2j})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$.
 $\Box$
@@
 $\Box$
-Proposition 16    Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class.
+Let $\Omega$ be a probability space with a probability measure $\mu$.
+Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$).
+Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$.
+Theorem 16    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep}q}$ are real-valued.
+Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
+Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$.
+For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
+$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$
+for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
+To show Theorem 16, we use the $\mathcal{A}$-valued Rademacher complexity defined by Hashimoto et al.
+(2023a).
+Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d.
+Rademacher variables, which take their values on $\{-1,1\}$.
+For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$.
+Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$.
+Similar to the case of the standard Rademacher complexity, an $\mathcal{A}$-valued version of the Rademacher complexity is defined as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$.
+Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
+Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity.
+In the following, we denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
+Proposition 20
+Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class.
 Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$.
@@
 where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued. $\Box$
-Theorem 15    Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep}q}$ are real-valued.
-Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
-Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$.
-For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
-$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$
-for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
-According to Hashimoto et al.
-[14], the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
+We obtain the following lemma directly by Lemma 4.2 in Mohri et al. (2012).
 Lemma 21
-Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq C\}$. Then, we have
-$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})\leq_{\mathcal{A}}\frac{B}{N}\bigg{(}\sum_{i=1}^{N}\|k(x_{i},x_{i})\|_{\mathcal{A}}\bigg{)}^{1/2}1_{\mathcal{A}}.$
-We obtain the following lemma directly by Lemma 4.2 in [31].
-Lemma 18
+Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class.
 Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function.
@@
 Lemma 22
-Let $g$ be the same map defined in Lemma 18. For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
+Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class.
+Let $g$ be the same map defined in Lemma 21.
+For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
 $\mathrm{E}[g(f(X),Y)](z)\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
+Here, $\mathrm{E}_{\sigma_{i}}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}]=\int_{\Omega}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}(\omega)\mathrm{d}\mu(\omega)$.
 Proof  With probability at least $1-\delta$, we have
@@
 $\quad=\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
-The inequality follows by Theorem 3.1 in [31]. $\Box$
-The following lemma completes the proof of Theorem 15.
+The inequality follows by Theorem 3.1 in Mohri et al. (2012). $\Box$
+According to Hashimoto et al.
+(2023a), the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
+Lemma 23
+Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Then, we have
+$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})\leq_{\mathcal{A}}\frac{B}{N}\bigg{(}\sum_{i=1}^{N}\|k(x_{i},x_{i})\|_{\mathcal{A}}\bigg{)}^{1/2}1_{\mathcal{A}}.$
+Combining these results with the following lemma completes the proof of Theorem 16.
 Lemma 24
 Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3.
-Then, for $x\in\mathcal{A}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
+Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
 Proof  We evaluate $\|k_{n}(x,x)\|_{\mathcal{A}}$. For $k_{n}^{\operatorname{poly},q}$, we have
-$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{\|}_{\operatorname{op}}$
-$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x)\|^{2q}_{\operatorname{op}}.$
+$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(x_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(x_{i})^{q}\bigg{\|}_{\operatorname{op}}$
+$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x_{i})\|^{2q}_{\operatorname{op}}.$
 The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
@@
 for any $z\in\mathbb{T}$.
-Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$.
-We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ [32, Chapter IX, Section 2].
-In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$. Thus, we have
+Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}^{n\times n}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$.
+We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ (Conway, 2007, Chapter IX, Section 2).
+In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$ and $\|v\|=1$. Thus, we have
 $\|R_{n}(x)\|_{\operatorname{op}}$ $=\|R_{n}(x)v\|=\|Q_{n}^{*}M_{x}Q_{n}v\|=\bigg{\|}Q_{n}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}\leq\bigg{\|}Q_{n+1}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}$
@@
 Appendix B Details of Remark 11
-To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 9 from the perspective of $q$. The following lemma is by Brandlini and Travaglini [25].
+To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 10 from the perspective of $q$.
+The following lemma is by Brandlini and Travaglini Brandolini & Travaglini (1997).
 Lemma 25
@@
 Proof  Let $u_{j,t}(r)=\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}$. Then, we have
-$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j$
-$=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
-$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j,$
-where $\hat{\cdot}$ is the Fourier transform, and the third equality is by the Poisson summation formula. $\Box$
-To understand the convrgence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$.
+$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}u_{j,t}(r)\mathrm{d}j$
+$=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
+$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
+$=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}nr\cdot m}\mathrm{d}r\mathrm{d}j,$
+where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula. $\Box$
+To understand the convergence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$.
 In the convergence discussed in Lemma 10, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$.
-On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 [25, Theorem 1].
+On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini & Travaglini, 1997, Theorem 1).
 Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@
 Hashimoto et al.
-[33, Proposition 4.5] showed an approximate representer theorem for RKHMs over von Neumann-algebras.
+(Hashimoto et al., 2023b, Proposition 4.5) showed an approximate representer theorem for RKHMs over von Neumann-algebras.
 Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
@@
 Proof  For $f_{1},f_{2}\in\mathcal{M}_{k}$, let $\theta_{f_{1},f_{2}}:\mathcal{M}_{k}\to\mathcal{M}_{k}$ defined as $\theta_{f_{1},f_{2}}(v)=f_{1}\left\langle f_{2},v\right\rangle_{k}$ for $v\in\mathcal{M}_{k}$.
-Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, and let $\mathcal{B}=\overline{\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$.
-In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see [12, p8]).
-Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), it has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ [23, Theorem I.4.8].
+Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, let $\mathcal{B}_{0}={\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$, and let $\mathcal{B}=\overline{\mathcal{B}_{0}}$.
+In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see (Lance, 1995, p8)).
+Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), $\mathcal{B}_{0}$ has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ (Davidson, 1996, Theorem I.4.8).
 For each $v\in\mathcal{N}_{k}$, we have
@@
 Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition. $\Box$
-Appendix D Experimental Details and Additional Results
+Appendix D Generalization to Other $C^{*}$-algebras
+Setting $R_{n}$ and $S_{n}$ for more general $C^{*}$-algebras has been investigated.
+Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
+Continuous functions on high-dimensional torus (Leimbach & van Suijlekom, 2024)
+[…]
+Continuous functions on the sphere (Rieffel, 2004)
+For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space.
+Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$.
+We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
+Appendix E Experimental Details and Additional Results
 We provide experimental details below. All the experiments in this paper were executed with Python 3.9 on an Intel(R) Core(TM) i9-10885H 2.4GHz processor with the Windows 10 operating system.
-Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order. (Average value of results of five different runs. The error bar represents the standard deviation.)
-D.1 Experiment with synthetic data
+Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task.
+(Average value of results of five different runs.
+The error bar represents the standard deviation.)
+E.1 Experiment with synthetic data
 We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$.
@@
 Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
-D.2 Experiment with MNIST
+We also show additional results regarding the deep approach in Figure 4.
+The result is the average $\pm$ the standard deviation of three independent runs.
+We set different values of $n$ with the same deep setting considered in Subsection 7.1.
+Figure 4: Test error of the regression task with deep approach with different $n$ and $L$.
+E.2 Experiment with MNIST
 We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.
+Figure 5 (a) shows the test error.
+We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$.
+In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0.
+Figure 5 (b) shows the result.
+The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$.
+Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
+Figure 5: (a) Test error of the image recovering task with different values of $n$ and the kernel $k=\hat{k}_{n}^{\operatorname{prod},q}$.
+(Box plot of results of five independent runs with different training and test data.) (b) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the image recovering task.
+(Average value of results of five different runs.
+The error bar represents the standard deviation.)

### v4

@@
 The noncommutativity induces interactions along the data function domain.
-We show that it is a governing factor leading to performance enhancement: we can balance the representation power and the model complexity.
-We also propose a deep learning perspective to increase the representation capacity of spectral truncation kernels.
-The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
+We show that the proposed kernels fill the gap between existing separable and commutative kernels.
+We also propose a deep learning perspective to obtain a more flexible framework.
+The flexibility of the proposed class of kernels allows us to go beyond previous separable and commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
+Keywords: Kernel methods, $C^{*}$-algebra, Positive definite kernel, Ridge regression
 1 Introduction
-Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf & Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017).
+Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf and Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017).
 They have been applied, for example, to ridge regression, principal component analysis, and support vector machine.
@@
 Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel.
-Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf & Smola, 2001; Thomas, 2008).
+Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf and Smola, 2001; Thomas, 2008).
 Standard positive definite kernels are scalar-valued, and are well-suited to learn scalar-valued functions.
@@
 A typical kernel is the separable kernel, which is defined by the product of a scalar-valued kernel and a positive semi-definite operator (Álvarez et al., 2012).
-Although applying the separable kernel is computationally efficient, it identifies dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables.
-Several attempts have been made to construct kernels that go beyond separable kernels.
-Huusari & Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement.
+Another typical kernel is the commutative kernel, which is defined only with the pointwise calculation of functions or vectors (Hashimoto et al., 2021).
+Although applying the separable and commutative kernels is computationally efficient, there is a crucial shortcoming for each kernel.
+Separable kernels identify dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables.
+The output is determined only by the global information of the input.
+On the other hand, commutative kernels only identify the pointwise (completely local) dependencies.
+Indeed, they are two extreme cases regarding the dependencies between input and output variables.
+Several attempts have been made to construct kernels that go beyond separable and commutative kernels.
+A typical nonseparable kernel is the transformable kernels, which is characterized by a map that can incorporate the information of input variables with the output variables.
+Huusari and Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement.
 Hashimoto et al.
@@
 Using this kernel, one can generalize the convolution and capture the effect of interactions of different Fourier components.
-However, if we need an $m$-dimensional vector-valued outputs with nonseparable kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general.
-Thus, with vvRKHSs, we cannot resolve the two issues at the same time.
-In this work, we address these two challenges by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021).
+However, if we need an $m$-dimensional vector-valued outputs with these kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general.
+Thus, with vvRKHSs, to go beyond separable and commutative kernels, the computational cost is significant.
+In this work, we address the two challenges of the computational cost and the choice of kernels by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021).
 RKHM is a generalization of RKHS by means of $C^{*}$-algebra.
@@
 By evaluating the values of the function-valued Gram matrix at $m$ different points, we obtain $m$ scalar-valued Gram matrices.
-This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels.
-To obtain kernels going beyond the separable kernels, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes & van Suijlekom, 2021).
-Our class of kernels is large and encompasses previous commutative kernel functions such as polynomial, product, and separable kernels.
+This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels such as transformable kernels.
+We summarize the difference between the proposed and existing kernels in Table 1.
+To obtain kernels going beyond the separable and commutative kernels with low computational cost, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes and van Suijlekom, 2021).
 The proposed kernels are parameterized by a natural number $n$ corresponding to the dimension of the truncated space.
@@
 Thus, $n$ describes the resolution of the discretization, and we call it the truncation parameter.
+Indeed, $n$ describes the noncommutativity of the kernel.
 We show that the proposed kernels converge to the commutative kernels as $n$ goes to infinity.
-In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process.
-The truncation parameter $n$ plays an important role from at least two perspectives.
-First, an important structure in typical commutative kernels is the product, and $n$ describes the noncommutativity of the product.
-For two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$.
-Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product.
-This situation corresponds to the case of $n=\infty$.
-On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$.
-This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$.
-This corresponds to the case $n<\infty$.
-Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity.
-We show that if $n$ is small, then the representation power is low, and the model complexity is small.
-On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large.
-The expectation loss for learning a function is controlled by the representation power and the complexity.
-By introducing the parameter $n$ and setting $n$ to balance them, we obtain higher performance compared to the typical commutative kernels.
-In addition, to resolve the issue of the interactions becoming weaker as $n$ becomes larger (i.e., representation power becomes higher), we propose a deep approach to obtain models with higher representation powers while keeping the interactions.
+On the other hand, if $n=1$, then the proposed kernels are separable kernel.
+Thus, we can control local and global dependencies through $n$.
+By introducing the parameter $n$ and setting $n$ to balance the information of gloal dependencies and that of the local dependencies, we obtain higher performance compared to the separable and commutative kernels.
+We also propose a deep model with the proposed kernels to make the framework more flexible and capture both local and global dependencies of the output on the input.
+In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), we have focused on how to interpolate observed data to obtain functions (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021).
+On the other hand, the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process.
 Our contributions are summarized as follows:
-  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that generalize typical kernels.
-The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which induces the interactions along the data function domain by introducing the noncommutativity into the learning process.
-(see Definition 6 and Subsection 3.1)
-  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels.
-The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity.
-(see Section 4)
-  • We propose a deep approach to gain the representation power while keeping the noncommutativity in the kernels. (see Section 6)
-Our framework allows us to go beyond the usual commutative kernels and to develop a new perspective on the significance of noncommutative structures in learning systems.
+  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that go beyond separable and commutative kernels with low computational cost. (see Definition 6)
+  • The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which adjusts the global and local interactions along the data function domain by introducing the noncommutativity into the learning process.
+(see Subsection 3.1)
+  • We propose a deep approach to make the framework more flexible (see Section 6).
+Table 1: Summary of the existing and the proposed kernels
+Type of kernels Computational cost (ridge regression) Extraction of local information Extraction of global information
+Separable $O(mN^{3})$$\times$ $\checkmark$
+Commutative $O(mN^{3})$checkmark $\times$
+Transformable $O(m^{3}N^{3})$checkmark $\checkmark$
+Proposed (with RKHM) $O(mn^{2}N^{2}+mN^{3})$checkmark $\checkmark$
 2 Preliminaries
+We review some mathematical notions and existing studies.
 2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
@@
 2.2 Spectral truncation on the torus
-The product in $C(\mathbb{T})$ is commutative.
-However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure.
-Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$.
+As we mentioned in Example 1, the product in $C(\mathbb{T})$ is commutative.
+However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a matrix, we can obtain a noncommutative product structure.
+Let $e_{j}$ be the Fourier function defined as $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$.
 Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square-integrable functions on $\mathbb{T}$.
@@
 where $F_{n}(t)=(1/n)\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution.
-By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ (Brandolini & Travaglini, 1997; van Suijlekom, 2021).
+If $n=1$, then $F_{n}(z)=1$, and as $n$ goes to infinity, $F_{n}$ goes to the delta function.
+More precisely, the following proposition holds (Brandolini and Travaglini, 1997; van Suijlekom, 2021), which implies that for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$.
 Proposition 5
 For each $z\in\mathbb{T}$, $x\ast F_{n}(z)\to x(z)$ as $n\to\infty$.
-A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini & Travaglini, 1997).
+A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini and Travaglini, 1997).
 Let $P$ be a polyhedron and $jP=\{jz\,\mid\,z\in P\}$ for $j\in\mathbb{Z}$.
@@
 $F_{n}^{q,P}(t)=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}.$ (1)
+2.3 Existing operator-valued kernels
+We review existing operator-valued kernels and discuss their advantages and shortcomings.
+We especially focus on the case $\mathcal{X}=\mathcal{A}^{d}$ and the application of kernel ridge regression in this paper.
+For more details about the operator-valued kernels, see, for example, Álvarez et al.
+(2012).
+Typical existing kernels are summarized as follows: Let $\tilde{k}:\tilde{\mathcal{X}}\times\tilde{\mathcal{X}}\to\mathbb{C}$ be a complex-valued positive definite kernel, $\mu$ be a finite positive measure on $\mathbb{T}^{2}$, and $S:\mathbb{T}\times\mathcal{X}\to\tilde{\mathcal{X}}$.
+Consider a kernel $k:\mathcal{X}\times\mathcal{X}\to\mathcal{B}(L^{2}(\mathbb{T}))$ defined by
+$\left\langle u,k(x,y)v\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}^{2}}u(z)\tilde{k}(S(z,x),S(t,y))v(t)\mathrm{d}\mu(t,z)$ (2)
+for $u,v\in L^{2}(\mathbb{T})$.
+In general, we have to construct a Gram matrix whose elements are operators.
+Typically, if we need an output function evaluated at $m$ points, then we discretize the operator on $O(m)$-dimensional space and obtain an $O(mN)$ by $O(mN)$ Gram matrix.
+Then, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m^{3})$.
+Therefore, the computational cost is significant.
+As a special case of the kernel (2), the following two kernels are efficient in the sense of the computational cost.
+Separable kernel
+If we set $\tilde{\mathcal{X}}=\mathcal{X}$ and $S(t,x)=x$, then the kernel $k$ defined in Eq.
+(2) is called a separable kernel.
+In other words, let $A\in\mathcal{B}(L^{2}(\mathbb{T}))$ be a Hermitian positive semi-definite operator defined as $\left\langle u,Av\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}}\overline{u(t)}v(z)\mathrm{d}\mu(z,t)$.
+The kernel $k:\mathcal{X}\times\mathcal{X}\to\mathcal{B}(L^{2}(\mathbb{T}))$ defined as
+$k(x,y)v(z)=\tilde{k}(x,y)Av(z)$
+is called a separable kernel.
+For separable kernels, the computation is reduced to the computation of the Gram matrix of $\tilde{k}$.
+Thus, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m)$, where $N$ is the sample size.
+However, as we can see from the definition, it identifies dependencies between input and output variables separately and cannot reflect local information of input variables to output variables.
+Indeed, with separable kernels, we cannot specify the relationship between the value of the output function at $z$ and the value $x(z)$ of the input $x$ at $z$ although they have strong connection in many cases.
+For example, if we have an input $[x_{1},\ldots,x_{d}]\in\mathcal{A}^{d}$ as $d$ time-series explanatory variables (i.e., $x_{1},\ldots,x_{d}$ are functions with respect to time) and try to obtain an output function as a response variable, the values of $x_{1}(z),\ldots,x_{d}(z)$ at time $z$ is strongly related to the value of the output at time $z$.
+In this case, the separable kernels are not suitable for extracting the relationship between $x_{1}(z),\ldots,x_{d}(z)$ and the values of the output at $z$.
+Commutative kernel
+[…]
+$k(x,y)v(z)=\tilde{k}(x(z),y(z))v(z).$
+We call this kernel the commutative kernel.
+For commutative kernels, the computation is reduced to the computation at each $z$.
+Thus, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m)$.
+However, as we can see from the definition, it only identifies completely local relationship between the input function and the output function.
+The value of the output function at $z$ is determined only with the value of the input function at $z$.
+In the same example as separable kernels, the values of $x_{1}(z),\ldots,x_{d}(z)$ at time $z$ is strongly related to the value of the output at time $z$, but may also be related to $y(z+t)$ for $t\in[-T,T]$ for a small number $T$.
+In this case, the commutative kernels are not suitable for extracting the relationship between $x_{1}(z),\ldots,x_{d}(z)$ and the values of the output around $z$, not only exactly at $z$.
+As we have seen, although the separable and commutative kernels are computationally efficient, they are two extreme cases regarding the description of the relationship between the input and output functions; the separable kernels identify dependencies between input and output variables separately (only with global information) and the commutative kernels only identify the commutative (completely local) dependencies.
+The following existing kernels fill a gap between the above two kernels, but computationally expensive.
+Transformable kernel
+If we set $\mu$ as the Lebesgue measure, then the kernel $k$ defined in Eq.
+(2) is called the transformable kernel.
+In this case, the kernels can identify the dependency between input and output variables through the map $S$.
+However, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points is $O(N^{3}m^{3})$ as we discussed at the beginning of this section.
+In addition, we have to determine the map $S$, but it is not easy to interpret in general.
+Combination of separable and transformable kernels
+Considering the sum of separable and transformable kernels and the product of these kernels have also been proposed. However, we have the same shortcomings of separable and transformable kernels.
 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
@@
 However, as we stated in Section 1, to go beyond the separable kernel and reach higher performance, the computational cost scales as $O(m^{3}N^{3})$ for $m$-dimensional outputs with $N$ samples.
-To resolve this issue, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs.
+(See Appendix 2.3.) To go beyond the separable and commutative kernels with lower computational cost, we consider function-valued ($\mathcal{A}$-valued) kernels and RKHMs.
 As we will discuss in the last part of Section 5, the computational cost depends on $mN^{3}$ if we use an $\mathcal{A}$-valued kernel.
-Hashimoto et al.
-(2021) proposed to use the $\mathcal{A}$-valued kernel $k$ where $k(\cdot,\cdot)(z)$ is a complex-valued positive definite kernel for all $z\in\mathbb{T}$.
-Combining it with typical existing kernels (Hashimoto et al., 2023a; Schölkopf & Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012), we obtain the following three kinds of $\mathcal{A}$-valued kernels.
+We first combine existing kernels (Hashimoto et al., 2021, 2023a; Schölkopf and Smola, 2001; Kadri et al., 2012; Álvarez et al., 2012) and obtain the following typical commutative $\mathcal{A}$-valued kernels.
+Here, we consider the case where the set $\mathcal{X}$ for data is $\mathcal{X}\subseteq\mathcal{A}^{d}$.
 Example 2
@@
   2. 2.
-    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$,
-  3. 3.
-    separable kernel $k^{\operatorname{sep},q}(x,y)=\tilde{k}(x,y)\prod_{j=1}^{q}(a_{j}^{*})^{q}a_{j}^{q}$.
-Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ and $\tilde{k}:\mathcal{A}^{d}\times\mathcal{A}^{d}\to\mathbb{C}$ are complex-valued continuous positive definite kernels.
-In addition, $a_{j}\in\mathcal{A}$ is the parameter of $k^{\operatorname{sep},q}$.
+    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$.
+Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ is a complex-valued continuous positive definite kernel.
 We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the typical kernels in Example 2.
@@
 For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1).
-Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$.
+Thus, as we also discussed in Subsection 2.3, the commutative kernels defined in Example 2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$.
 If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship.
@@
 $k_{n}^{\operatorname{poly},q}(x,y)$ $=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)},$
-$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)},$
-$k_{n}^{\operatorname{sep},q}(x,y)$ $=\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j})\bigg{)}.$
+$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)}.$
 Here, $A^{*}$ for a matrix $A$ is the adjoint, and we denote by $\tilde{k}_{i,j}(x,y)$ the map $z\mapsto\tilde{k}_{i,j}(x(z),y(z))$.
@@
 […]
+3.1 Connection between the proposed kernel and existing kernels
+The proposed function-valued kernels combined with the framework of RKHMs fill the gap between separable and commutative kernels with a computational cost lower than that of transformable kernels.
+The proposed kernels are separable if $n=1$ and commutative if $n=\infty$.
+Indeed, we have the following observation:
+The case of $n=1$
+The proposed kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ are equivalent to separable kernels:
+$k_{1}^{\operatorname{poly},q}(x,y)(z)$ $=\sum_{i=1}^{d}\alpha_{i}\bigg{(}\int_{\mathbb{T}}\overline{x_{i}(t)}\mathrm{d}t\bigg{)}^{q}\bigg{(}\int_{\mathbb{T}}y_{i}(t)\mathrm{d}t\bigg{)}^{q},$
+$k_{1}^{\operatorname{prod},q}(x,y)(z)$ $=\prod_{j=1}^{q}\int_{\mathbb{T}}\overline{\tilde{k}_{1,j}(x(t),y(t))}\mathrm{d}t\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),y(t))\mathrm{d}t.$
+We can see that the output of the kernel does not depend on the variable $z$.
+For example, the output function with an input $x$ of the kernel ridge regression is in the form of $\sum_{i=1}^{N}k(x,x_{i})(z)c_{i}(z)$ for given samples $x_{1},\ldots,x_{N}\in\mathcal{X}$ and some $c_{1},\ldots,c_{N}\in C(\mathbb{T})$ (See Hashimoto et al.
+(2023a) or Section 5).
+Thus, the kernels identify dependencies between input and output functions separately.
+We can only capture global information of the input $x$ for obtaining the output.
+The case of $n=\infty$
+In the next section (Theorem 10), we will show that the proposed kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ converge to commutative kernels $k^{\operatorname{poly},q}$ and $k^{\operatorname{prod},q}$ as $n$ goes to infinity.
+Thus, the proposed kernels are equivalent to commutative kernels:
+$k_{\infty}^{\operatorname{poly},q}(x,y)(z)$ $=\sum_{i=1}^{d}\alpha_{i}(\overline{x_{i}(z)}y_{i}(z))^{q},$
+$k_{\infty}^{\operatorname{prod},q}(x,y)(z)$ $=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z)).$
+The value of the output of the kernels at $z$ are determined only by $x(z)$ and $y(z)$. The kernels identify completely local dependencies between input and output functions.
+The case of $1<n<\infty$
+If $n$ is small, then we focus more on global information, and if $n$ is large, we focus more on local information. Indeed, as we discussed in Eq. (3), we have
+[…]
+where we assume $d=1$ for simplify the notation.
+The proposed kernels are described by the convolution of the input function and the Féjer kernel $F_{n}^{2q,P}$.
+If $n=1$, then $F_{n}^{2q,P}(z)=1$, and as $n$ goes to infinity, $F_{n}^{2q,P}$ goes to the delta function.
+This means that if $n$ is small, then the convolution with $F_{n}^{2q,P}$ extract global information more than local information.
+On the other hand, if $n$ is large, then the convolution with $F_{n}^{2q,P}$ extract local information more than global information.
+We illustrate how $F_{n}^{2,P}$ changes along $n$ in Figure 2.
+Figure 2: Féjer kernel $F_{n}^{2,P}$ for $n=5,10,15$
+Remark 8
+We can determine optimal $n$ in the sense of the dependencies by observing the Féjer function $F_{n}^{2q,P}$.
+Since the proposed kernel is defined by the convolution of the input function with the Féjer kernel, the volume of the region where the value of the Féjer kernel is sufficiently large corresponds to the range of local dependencies.
+Thus, if we have a information of the local dependencies, then we can choose $n$ based on the values of $F_{n}^{2q,P}$.
+We also discuss a generalization of the proposed kernel to kernels with a continuous value parameter, instead of the discrete value parameter $n$, in Subsection 9.2.
+This generalization enables us to use an optimization method to learn a best possible parameter.
+Remark 9
+The proposed kernels are composed of complex-valued kernels $\tilde{k}_{i,j}$ in Definition 6.
+We can choose any kernel for $\tilde{k}_{i,j}$, and the properties of the proposed kernels depend also on that choice.
+If we choose a kernel with parameters as $\tilde{k}_{i,j}$ (such as a weighted sum of multiple kernels), then by optimizing the parameters, we can obtain a better kernel for given data or tasks.
+In addition, for function-valued kernels, in the same manner as the complex-valued kernels, the weighted sum of positive definite kernels and product of positive definite kernels are also a positive definite kernel.
+Thus, we can consider multiple proposed kernels and combine them with weight parameters.
+In that case, we can also optimize the parameters to obtain a better kernel for given data or tasks.
+Computational cost
+As we will discuss in the last paragraph of Section 5, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points with the proposed kernel $k_{n}^{\operatorname{poly},q}$ or $k_{n}^{\operatorname{prod},q}$ is $O((q+m)n^{2}N^{2}+mN^{3})$.
+Thus, if $(q+m)n^{2}<m^{3}N$, then the proposed kernels are more computationally efficient than nonseparable and noncommutative kernels, such as the transformable kernels.
 3.2 Convergence and interactions
-We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2 as $n$ goes to infinity, respectively.
-In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of typical kernels.
-In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$.
+We show theoretical results that support the discussion in Subsection 3.1.
+We first show that $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$ and $k^{\operatorname{prod},q}$ defined in Example 2 as $n$ goes to infinity, respectively.
 In the following, all the proofs of theoretical statements are documented in Appendix A.
@@
 For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$.
-If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
 As we can see in Eq.
@@
 Then, for any $g\in C(\mathbb{T}^{q})$ and any $z\in\mathbb{T}^{q}$, $g\ast F_{n}^{q,P}(z)\to g(z)$ as $n\to\infty$.
-Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 9 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq.
+Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 11 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq.
 (3).
-Lemma 10 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini & Travaglini (1997).
+Lemma 12 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is proved by Brandolini and Travaglini (1997).
 We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq.
-(2), and apply Lemmas 9 and 10 to show Theorem 8.
+(3), and apply Lemmas 11 and 12 to show Theorem 10.
 See Appendix A for more details.
-As seen in Theorem 8, the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows.
-On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel.
-Regarding the relationship between $q$ and the interactions, we have the following remark.
-Remark 11
-We expect that the convergence in Theorem 8 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.
-This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$.
-See Appendix B for more details.
+Theorem 10 implies that the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows.
+This is because as $n$ goes to infinity, $F_{n}^{2q,P}$ goes to the delta function and by taking the convolution with the input function, it focuses more on local relationships between input and output functions as $n$ grows.
+Regarding the relationship between $q$ and the interactions see Appendix B.
 3.3 Positive definiteness
@@
 Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product.
-Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
 Proposition 13
-The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
+The kernel $k_{n}^{\operatorname{poly},q}$ is positive definite.
 As for $k_{n}^{\operatorname{prod},q}(x,y)$, it depends on $x$ and $y$ through $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products.
@@
 Lemma 15
-We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
+The Féjer kernel $F_{n}^{q,P}$ is bounded as $|F_{n}^{q,P}(z)|\leq n^{q}$.
 Remark 16
@@
 In addition, considering non-positive kernels has also been investigated (Ong et al., 2004; Canu et al., 2005).
-Indeed, in practical computations in Subsection 7, the algorithms can work even if we set $\beta_{n}<n^{2q}$.
+Indeed, in practical computations in Section 8, the algorithms can work even if we set $\beta_{n}<n^{2q}$.
 Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@
 We now focus on what impact the truncation parameter $n$ has on generalization.
-We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity associated with the proposed kernels.
 According to Lemma 4.2 in Mohri et al.
@@
 Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous.
-We derive the following generalization bound for the kernels defined in Section 3.
+We derive the following generalization bound for the kernels defined in Section 3 based on Mohri et al.
+(2012); Hashimoto et al.
+(2023a) Here, to adapt to the generalization bound analysis, we assume the kernels are real-valued.
 Theorem 17
-Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are real-valued.
-Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
+Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$ are real-valued.
+Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
 Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$.
-For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
+For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
 $\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$ (5)
@@
 If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ in $\mathcal{A}$ is larger, the outputs of ${k_{n}}$ can represent more functions, which enables us to describe the dependency of outputs on inputs more.
-Indeed, if $n=1$, then the output of $k_{n}$ is always a constant function, where the situation is the same as the existing separable kernel discussed in Section 1.
 Thus, the empirical loss (the first term) can become small if $n$ is large.
-On the other hand, according to Theorem 16, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger.
-An advantage of introducing the parameter $n$ to construct the new kernels based on the typical kernels in Example 2 is that we can control the empirical loss and the complexity through $n$.
-By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$, the left-hand side of Eq.
-(3), small.
-5 Application to Kernel Ridge Regression
+On the other hand, the complexity of $\mathcal{M}_{k_{n}}$ (the second term of the right-hand side of (5)) is larger if $n$ is larger.
+5 Kernel Ridge Regression with the Spectral Truncation Kernels
 We illustrate the effect of the proposed kernels by applying them to kernel ridge regression.
 Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples.
-We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$.
+We consider the case where we need the values of the output function evaluated at different $m$ points.
+We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$ or $k=k_{n}^{\operatorname{prod},q}$.
 Consider the typical minimization problem for regression, which is also considered by Hashimoto et al.
@@
 $\quad=\min_{c_{j}\in\mathcal{A}}(\mathbf{c}^{*}\mathbf{G}^{2}\mathbf{c}-\mathbf{c}^{*}\mathbf{G}\mathbf{y}-\mathbf{y}^{*}\mathbf{G}\mathbf{c}+\lambda\mathbf{c}^{*}\mathbf{G}\mathbf{c}),$ (7)
-where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$.
+where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, and $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$.
 Then, the solution $\mathbf{c}$ of the problem (7) satisfies $\mathbf{y}=(\mathbf{G}+\lambda I)\mathbf{c}$, which means $\mathbf{y}(z)=(\mathbf{G}(z)+\lambda I)\mathbf{c}(z)$.
@@
 The situation is different if we use vvRKHSs.
-For vvRKHSs, to obtain $m$ values as an output, we apply a $\mathbb{C}^{m\times m}$-valued kernel.
-If the kernel is nonseparable, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$.
-We can reduce the cost with respect to $N$, e.g., by Nyström method (Drineas & Mahoney, 2005).
+For vvRKHSs, to obtain $m$ values as an output, we typically apply a $\mathbb{C}^{m\times m}$-valued kernel.
+Then, the Gram matrix should be $mN$ by $mN$.
+Specifically, if the kernel is transformable, the computational cost is $O(m^{3}N^{3})$.
+Thus, if $(q+m)n^{2}<m^{3}N$, then the proposed kernels are more computationally efficient than the transformable kernels.
+For the proposed kernel, we can reduce the cost with respect to $N$, e.g., by applying Nyström method (Drineas and Mahoney, 2005) to each $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$.
 Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel.
+Generalization of the proposed kernel with the Fejér kernel to a kernel with another integral kernel, which we will discuss in Subsection 9.2 is also a way to reduce computational cost.
 However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
 6 Deep model with the spectral truncation kernels
-As we discussed in Section 4, there is a tradeoff between the representation power and the model complexity.
-In fact, there are two directions of the representation power: one is about the dependency of the outputs on the inputs discussed in Section 4, and the other is about the interactions along the data function domain.
-One shortcoming of the spectral truncation kernel is that to obtain a model with a high representation power in the sense of the dependency of the outputs on the inputs, the truncation parameter $n$ have to be large, which results in making the interactions along the data function domain weak.
-To gain the two types of representation powers at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n}^{1},\ldots,k_{n}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
-$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (6)
-We consider the case where $c_{i}^{j}$ in Eq. (6) is represented as a finite sum of functions in $\mathcal{A}$. In the following case, the representation power of this model grows exponentially.
+As we discussed in Subsection 3.1, the truncation parameter $n$ controls the global and local dependencies of the output on the input.
+To make the framework more flexible and capture both global and local dependencies at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n_{1}}^{1},\ldots,k_{n_{L}}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
+$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n_{j}}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (8)
+If we set a different value of $n_{j}$ for each layer $j$, then we can capture different level of global and local dependencies for each layer, which enables us to capture both global and local dependencies.
+To learn the deep model, we need to represent $c_{i}^{j}$ in Eq.
+(8) as a finite sum of functions in $\mathcal{A}$.
+The following proposition provides an appropriate way to parameterize $c_{i}^{j}$ in the sense that the representation power of the model grows exponentially.
+For simplicity, we consider the case where $n_{j}=n$ for $j=1,\ldots,L$ in the following proposition.
+However, we can derive the same result for the case of $n_{j}\neq n_{l}$ for $j\neq l$.
 Proposition 18
@@
 Remark 19
-The exponential growth of the representation power is also observed for neural networks (Hanin & Rolnick, 2019).
+The exponential growth of the representation power is also observed for neural networks (Hanin and Rolnick, 2019).
 Deep learning with kernels has also been proposed (Laforgue et al., 2019; Hasimoto et al., 2023).
@@
 Remark 20
-The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k^{j}$.
+The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k_{n_{j}}^{j}$.
 Since the Rademacher complexity bound depends on the norm of the kernel, it can also grow exponentially with respect to $L$.
@@
 Investigating how we can take advantage of the deep setting in the sense of the generalization is future work.
+7 Applications: Extracting Local and Global Dependencies of functions
+There are many potential applications of the spectral truncation kernels. We list two examples here.
+Time-series data analysis
+We can regard a time-series as a function on a time space.
+In many cases, a state at a certain time $z$ is influenced strongly by another state at the same time $z$, but also by the state around the time $z$.
+Since commutative kernels focus only on local information, we cannot describe these two states with commutative kernels.
+On the other hand, since separable kernels focus only on global information, we cannot describe the relationship of these two states at each time $z$.
+By applying the proposed kernels, we can extract global information, but also can focus on local information.
+Operator learning
+In the framework of operator learning, we obtain a solution of a partial differential equation as an output from an input function (such as initial condition or parameter of the equation).
+Thus, we construct a model where both of the input and output are functions.
+Applying kernel methods to operator learning has been proposed (Batlle et al., 2024).
+We can construct the model by solving a kernel ridge regression task.
+Another well-known operator learning method is neural operator.
+In the framework of neural operators, we apply integral operators to extract global information and apply local linear operators and local activation functions to extract local information.
+The proposed kernel enables us to do similar procedures for the operator learning with kernels.
+By considering the product of multiple proposed kernels with different values of $n$ or deep model with the proposed kernels with different values of $n$, we can extract both global and local information; we can extract global information using the kernel with small $n$ and extract local information using the one with large $n$ in the model.
+We will numerically confirm the applicability of the spectral truncation kernels to operator learning in Subsection 8.3.
 8 Numerical results
@@
 8.1 Experiment with synthetic data
-We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compared them to the commutative kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$).
+We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$ and $\hat{k}_{n}^{\operatorname{prod},q}$ with different values of $n$ to observe that the truncation parameter $n$ controls global and local dependencies of the output on the input, as discussed in Subsection 3.1.
 We considered a regression task with synthetic data.
@@
 We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$.
-For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1\cdot 2\pi^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=2\pi\mathrm{e}^{\sin z}$.
-We estimated $f$ using kernel ridge regression.
+We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$.
+We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $(1/N)\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.
+For $x^{i}$, $y^{i}$, and $\tilde{x}^{i}$, we discretized each function with 30 equally spaced points on $\mathbb{T}$.
 We applied the cross-validation grid search to find an optimal regularization parameter $\lambda$.
-Figure 2 (a-c) illustrates the test error.
+Figure 3 (a,b) illustrates the test error.
 We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$.
 It becomes large if $n$ is larger or smaller than $32$.
-This is because when $n=32$, the empirical error and the complexity in Eq.
-3 are balanced, and the best possible expected error is obtained.
+This is because when $n=32$, the information of the global and that of the local dependencies are balanced, and the best possible error is obtained.
 We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the typical commutative kernel ($n=\infty$).
-We have similar results for the other two kernels.
+We have similar results for ${k}_{n}^{\operatorname{prod},q}$.
 Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$.
@@
 However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case.
-See Appendix E.1 for more details.
-We also compared the performance and the computational time of RKHMs associated with the proposed kernels to a vvRKHS with the same setting as above.
-For the vvRKHS, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$.
-This nonseparable kernel is proposed by Lim et al.
+To investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0.
+Figure 3 (c) shows the result.
+The index $i$ is determined in the descendant order with respect to the value of $\lambda_{i}$.
+Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
+We also compared the performance and computational time with the proposed kernels to the existing operator-valued kernel with the same setting as above to show the advantage of the proposed kernel in the sense of the computational efficiency and performance.
+For the existing kernel, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$.
+This kernel is the combination of the separable and transformable kernels and proposed by Lim et al.
 (2015).
 Table 2 shows the result.
-We can see that the RKHMs outperform the vvRKHS, and the computational time for the RKHMs are smaller than that for the vvRKHS.
-Note that we have already confirmed the proposed kernels outperform the existing separable kernel $k_{\infty}^{\operatorname{sep},q}$ in the above experiment (See Figure 2 (c)).
-Finally, we observed the behavior of the deep model.
+We can see that the proposed kernels outperform the existing typical nonseparable kernel, and the computational time with the proposed kernel is smaller than that with the existing kernel.
+Finally, we observed the behavior of the deep model to show its high representation power.
 We generated data in the same manner as above, but we set $N=300$.
 We used the kernel ${k}_{n}^{\operatorname{prod},q}$ and set $L=1,2,3$, $\tau_{1}=\sqrt{2}$, $\tau_{2}=\sqrt{3}$, and $\tau_{3}=\sqrt{5}/2$.
-We set $n$ so that the numbers of parameters are the same for all values of $L$.
+We set $n_{j}=n$ for $j=1,\ldots,L$ for simplicity and set $n$ so that the numbers of parameters are the same for all values of $L$.
 We used the loss function $\|\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{k}\|_{L^{2}(\mathbb{T})}+0.1\||f|_{k}\|_{L^{2}(\mathbb{T})}$ for the deep model $f$ defined in Eq.
 (8).
-Figure 2 (d) shows the result.
+Figure 4 shows the result.
 We can see that as the number of layers increases, we can obtain a higher performance model even with the same number of parameters.
@@
 (b) $k=\hat{k}_{n}^{\operatorname{prod},q}$
-(c) $k=k_{n}^{\operatorname{sep},q}$
-(d) Deep approach
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
-(e) Output images
-Figure 2: (a-c) Test error of the regression task with different values of $n$.
-(Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the regression task with deep approach with different $n$ and $L$.
-The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases.
-(e) Output images of the image recovering task with different values of $n$.
+(c) Eigenvalues
+Figure 3: (a,b) Test error of the regression task with different values of $n$.
+(Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (c) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task.
+(Average value of results of five different runs.
+The error bar represents the standard deviation.)
 Table 2: Test error and CPU time of the regression task with RKHM and vvRKHS
@@
 RKHM, $k=k_{n}^{\operatorname{prod},q}$ ($n=9$) 0.0113$\pm$1.63×$\times 10^{-5}$ 149.3$\pm$2.392
-RKHM, $k=k_{n}^{\operatorname{sep},q}$ ($n=10$) 0.0385$\pm$5.04388×$\times 10^{-5}$ 22.79$\pm$ 0.5742
-vvRKHS 0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
+vvRKHS (combination of transformable
+and separable kernels)
+0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
+Figure 4: Test error of the regression task with deep approach with different $n$ and $L$.
+The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases.
+(Average value $\pm$ the standard deviation of three independent runs.)
+Figure 5: (a) Test error of the image recovering task with different values of $n$ and the kernel $k=\hat{k}_{n}^{\operatorname{prod},q}$.
+(Box plot of results of five independent runs with different training and test data.) (b) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the image recovering task.
+(Average value of results of five different runs.
+The error bar represents the standard deviation.)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
+Figure 6: Output images of the image recovering task with different values of $n$.
 8.2 Experiment with MNIST
-We consider an image recovering task with MNIST (LeCun et al., 1998).
+To check the performance of the proposed kernels for practical applications, we consider an image recovering task with MNIST (LeCun et al., 1998).
 For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00.
@@
 We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$.
-Figure 2 (e) shows the output images with different values of $n$.
+We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $(1/N)\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.
+Figure 5 (a) shows the test error.
+We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$.
+In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0.
+Figure 5 (b) shows the result.
+The index $i$ is determined in the descendant order with respect to the value of $\lambda_{i}$.
+Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
+Furthermore, Figure 6 shows the output images with different values of $n$.
 When $n=\infty$, i.e., the commutative kernel, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image.
 When $n=70$, we can recover the image the clearest.
-8 Conclusion and Limitation
-In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation.
+8.3 Experiment for operator learning
+To show the availability of the proposed kernels to operator learning, we conducted an experiment. Consider the following Burgers’ equation on $(0,1]\times(0,1]$ for $w(x,t)$.
+$\frac{\partial w}{\partial t}+w\frac{\partial w}{\partial x}$ $=\nu\frac{\partial^{2}w}{\partial x^{2}}\qquad x\in(0,1)\times(0,1],$
+$w(x,0)$ $=u(x)\qquad x\in(0,1),$
+where $\nu=0.1$.
+We learned the operator mapping the initial condition $u$ to $w(\cdot,1)$, the solution at time $t=1$, by kernel ridge regression explained in Section 5.
+The training data and test data are generated by sampling the initial condition $u$ from the Gaussian distribution with mean 0 and the covariance $625(-\Delta+25I)^{-2}$.
+We used equally spaced $128$ points for computing the integral related to the input functions.
+The number of training samples is 1000 and that of test samples is 200.
+The same problem is also considered by Batlle et al.
+(2024).
+We used the same $\hat{k}_{n}^{\operatorname{prod},q}$ kernel as Subsection 8.1 with $n=5$.
+We set the regularization parameter $\lambda$ as $0.01$.
+Figure 7 shows the true solution $w(\cdot,1)$ and the predicted function $v$ by using the proposed kernel for the same test input function $u$ (initial condition).
+Figure 8 shows the error $v-w(\cdot,1)$ of these two functions.
+The mean value of the pointwise error $(1/128)\sum_{i=1}^{128}|v(x_{i})-w(x_{i},1)|$ over all the test samples is $0.00157\pm 7.079\times 1e^{-6}$ (average $\pm$ standard deviation over five independent runs).
+We can see that the proposed kernel combined with kernel ridge regression properly predict the solution of the problem, and compared to the results in Subsection 4.3.1 by Batlle et al.
+(2024), the proposed kernel outperforms these existing results.
+Figure 7: The true solution $w(\cdot,1)$ and the predicted function $v$ of the Burgers’ equation.
+Figure 8: Pointwise error $v-w(\cdot,1)$ of the true solution $w(\cdot,1)$ and the predicted function $v$ of the Burgers’ equation.
+9 Generalization of the Spectral Truncation Kernels
+The proposed spectral truncation kernels have potential possibilities of constructing kernels for more general situations.
+9.1 Generalization to other $C^{*}$-algebras
+Setting $R_{n}$ and $S_{n}$ for more general $C^{*}$-algebras has been investigated.
+Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
+Continuous functions on high-dimensional torus (Leimbach and van Suijlekom, 2024)
+[…]
+Continuous functions on the sphere (Rieffel, 2004)
+For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space.
+Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$.
+We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
+9.2 Generalization using other integral kernels
+We can generalize the spectral truncation kernel to a kernel on a measure space $\mathcal{X}$ by replacing the function $F_{n}^{2q,P}$ in Eq.
+(4) with another function $G:\Theta\times\mathcal{X}^{2q}$ that satisfies $G(\theta,z)\geq 0$ for any $\theta\in\Theta$ and $z\in\mathcal{X}^{2q}$.
+Here, $\Theta$ is the space of parameters.
+For example, we can set $\Theta=\mathbb{R}_{+}$, the set of positive real values, $\mathcal{X}=\mathbb{R}^{d}$, and $G(\sigma,z)=\mathrm{e}^{-\sigma\|z\|^{2}}$.
+Then, similar to the spectral truncation kernel, the convolution with $G(\sigma,z-\cdot)$ captures global information if $\sigma$ is small and captures local information if $\sigma$ is large.
+In this case, since the parameter space $\Theta$ is continuous, we can use an optimization method to learn a best possible parameter $\sigma$ from data.
+10 Conclusion and Discussion
+In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation, which fills the gap between existing separable and commutative kernels.
 The proposed kernels with the framework of RKHMs resolve two shortcomings of the framework of vvRKHSs for vector- or function-valued outputs at the same time: computational cost and choice of kernels.
 By considering function-valued positive definite kernels, not operator-valued ones, we can alleviate the computational cost.
-At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain.
-We also showed that we can control the tradeoff between the representation power and the model complexity associated with the proposed kernels.
-In addition, we proposed a deep approach to obtain models with higher representation powers.
+At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain, which enables us to control the global and local dependencies of the output on the input.
+In addition, we proposed a deep approach to obtain models with more flexibility.
 In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus.
 Although this setting includes many important examples, e.g., periodic time-series data, continuous functions defined on a compact set in $\mathbb{R}$, finite-dimensional vectors, investigations for more general $C^{*}$-algebras allow us to apply the proposed kernel more general settings.
-As we discuss in Appendix D, we can generalize this setting and define positive definite kernels for more general inputs and outputs.
-However, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
+As we discuss in Section 9, we can generalize this setting and define positive definite kernels for more general inputs and outputs and for more flexible frameworks.
+In addition, as we discuss in Section 7, there are many potential applications of the proposed kernels.
+Theoretical and empirical investigations for these generalized settings and applications remain to be investigated in future work.
 Acknowledgements
@@
 Theorem 10    For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$.
-If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
-Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$.
-The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are similar.
-We have
+Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$. The proof for $k_{n}^{\operatorname{prod},q}(x,y)$ is similar. We have
 $k_{n}^{\operatorname{poly},q}(x,y)(z)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}(z)$
@@
 Since $P$ is a convex polyhedron, $g_{i}\ast F_{n}^{2q,P}(z)\to g_{i}(z)$ as $n\to\infty$ by Lemma 12.
-For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable (van Suijlekom, 2021, Lemma 10).
-$\Box$
 […]
@@
 =$=\bigg{\{}r\in\mathbb{R}^{2q}\,\bigg{|}\,\bigg{|}\sum_{i=l}^{k}r_{i}\bigg{|}\leq 1,\ l\leq k\bigg{\}},$
-which completes the proof of the lemma. $\Box$
-Proposition 12    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
+which completes the proof of the lemma.
+Proposition 13    The kernel $k_{n}^{\operatorname{poly},q}$ is positive definite.
 Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
@@
 $=\sum_{i=1}^{d}\alpha_{i}\sum_{m=1}^{n}\bigg{|}\sum_{j=1}^{N}\sum_{j^{\prime}=1}^{n}d_{j}(z)\big{(}(R_{n}(x_{i,j}))^{q}\big{)}_{m,j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}\geq 0$
-for $z\in\mathbb{T}$. In addition, let $\tilde{x}_{n,j}=S_{n}(R_{n}(x_{j}))$. Then, we have
-$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{sep},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
-$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{2q-1}=1}^{n}$
-$\big{(}(R_{n}(a)^{*})\big{)}_{j^{\prime},m_{1}}\big{(}(R_{n}(a)^{*})\big{)}_{m_{1},m_{2}}\cdots\big{(}(R_{n}(a)^{*})\big{)}_{m_{q-1},m_{q}}\big{(}R_{n}(a)\big{)}_{m_{q},m_{q+1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{2q-1},l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}$
-$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)$
-$\qquad\sum_{m_{q}=1}^{n}\bigg{|}\sum_{j^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{q-1}=1}^{n}\big{(}R_{n}(a)\big{)}_{m_{q},m_{1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{q-2},m_{q-1}}\big{(}R_{n}(a)\big{)}_{m_{q-1},j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}$
-$\geq 0.$
-Thus, $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite. $\Box$
+for $z\in\mathbb{T}$. Thus, $k_{n}^{\operatorname{poly},q}$ is positive definite.
 Proposition 14    Let $\beta_{n}\geq-\min_{z\in\mathbb{T}^{q}}F_{n}^{2q,P}(z)$. Then, $\hat{k}_{n}^{\operatorname{prod},q}$ defined below is positive definite.
@@
 The last inequality is derived since the map $([x_{1},\ldots,x_{2q}],[y_{1},\ldots,y_{2q}])\mapsto\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{2j-1},y_{2j-1})}\tilde{k}_{2,j}(x_{2j},y_{2j})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$.
-$\Box$
 Lemma 15    We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
@@
 $|F_{n}^{q,P}(z)|\leq\frac{1}{n}n^{q+1}=n^{q}.$
-$\Box$
 Let $\Omega$ be a probability space with a probability measure $\mu$.
@@
 Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$.
-Theorem 16    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep}q}$ are real-valued.
-Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
+Theorem 17    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$ are real-valued.
+Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$ and $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$.
 Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$.
-For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
+For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
 $\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$
@@
 $=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}\overline{f(x_{i})}\sigma_{i}\bigg{|}\bigg{]}=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}f(x_{i})\sigma_{i}\bigg{|}\bigg{]}=\hat{R}_{N}(\mathbf{x},\mathcal{F}(z)),$
-where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued. $\Box$
+where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued.
 We obtain the following lemma directly by Lemma 4.2 in Mohri et al. (2012).
@@
 Here, $\mathrm{E}_{\sigma_{i}}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}]=\int_{\Omega}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}(\omega)\mathrm{d}\mu(\omega)$.
-Proof  With probability at least $1-\delta$, we have
+Proof  Note that we have $g(f(X_{i})(z),Y_{i}(z))=g(f(X_{i}),Y_{i})(z)$. Therefore, with probability at least $1-\delta$, we have
 $\mathrm{E}[g(f(X),Y)](z)=\mathrm{E}[g(f(X)(z),Y(z))]$
-$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}$
-$\quad=\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
-The inequality follows by Theorem 3.1 in Mohri et al. (2012). $\Box$
+$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
+The inequality follows by Theorem 3.1 in Mohri et al. (2012).
 According to Hashimoto et al.
@@
 Lemma 25
-Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3.
-Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
+Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ defined in Section 3. Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
 Proof  We evaluate $\|k_{n}(x,x)\|_{\mathcal{A}}$. For $k_{n}^{\operatorname{poly},q}$, we have
@@
 $\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x_{i})\|^{2q}_{\operatorname{op}}.$
-The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
+The calculation for $\hat{k}_{n}^{\operatorname{prod},q}$ is similar. The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
 $|S_{n}(A)(z)|$ $=\bigg{|}\frac{1}{n}\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}\bigg{|}$
@@
 $\leq\|Q_{n+1}^{*}M_{x}Q_{n+1}\|_{\operatorname{op}}=\|R_{n+1}(x)\|_{\operatorname{op}},$
-where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function. $\Box$
-Appendix B Details of Remark 11
+where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function.
+Appendix B Relationship between the parameter $q$ and the interactions
+Theorem 10 implies that the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows.
+On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel.
+Regarding the relationship between $q$ and the interactions, we have the following remark.
+Remark 26
+We expect that the convergence in Theorem 10 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.
+This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$.
 To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 12 from the perspective of $q$.
-The following lemma is by Brandlini and Travaglini Brandolini & Travaglini (1997).
+We have the following lemma, which is also mentioned by Brandolini and Travaglini (1997).
 Lemma 27
@@
 $=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}nr\cdot m}\mathrm{d}r\mathrm{d}j,$
-where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula. $\Box$
+where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula.
 To understand the convergence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$.
 In the convergence discussed in Lemma 12, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$.
-On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini & Travaglini, 1997, Theorem 1).
+On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini and Travaglini, 1997, Theorem 1).
 Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@
 $n^{q}\int_{t\in\mathbb{T}^{q}}\int_{0}^{1}\int_{jP}\mathrm{e}^{\mathrm{i}nr\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}j\mathrm{d}t-g(z)=\int_{0}^{1}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j.$
-Let ${h}^{P}_{s}(z)=\int_{t\in\mathbb{T}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$.
 Then, $h^{P}_{0}(z)=0$, and since $\lim_{s\to\infty}h^{P}_{s}(z)=0$, there exists $D(z)>0$ such that $|h^{P}_{s}(z)|\leq D(z)$ for any $s>0$.
@@
 By setting $C(z)=\max\{\tilde{C}(z),|x_{2}(z_{2})|\}$, we obtain the result.
-$\Box$
-The left hand side is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$.
+The left-hand side of Eq.
+(9) is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$.
 The first term in the right hand side of Eq.
 (9) is described by the convergence of $h_{jn}^{[-1,1]}(z)$, which corresponds to the one-dimensional ($(q-1)$-dimensional) Fejér kernel.
-Since we have an additional term $1/2|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$.
+Since we have an additional term $(1/2)|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$.
 Based on this observation, we expect that the convergence in Theorem 10 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.
@@
 (Hashimoto et al., 2023b, Proposition 4.5) showed an approximate representer theorem for RKHMs over von Neumann-algebras.
-Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
+Here, von Neumann-algebras are $C^{*}$-algebras that is closed with respect to the strong operator topology.
+Since $C(\mathbb{T})$ is not a von Neumann-algebra, for the arguments in Section 5, we generalize the theorem to that for general $C^{*}$-algebras as follows.
 Proposition 29
@@
 Thus, $b_{i}-b_{i}^{2}=c^{2}-c^{4}=c(1-c^{2})c=c^{*}(1-b_{i})c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$.
-Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}$.
+Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}^{2}$.
 As a result, we obtain
@@
 $\leq_{\mathcal{A}}g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N})).$
-Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition. $\Box$
-Appendix D Generalization to Other $C^{*}$-algebras
-Setting $R_{n}$ and $S_{n}$ for more general $C^{*}$-algebras has been investigated.
-Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
-Continuous functions on high-dimensional torus (Leimbach & van Suijlekom, 2024)
-[…]
-Continuous functions on the sphere (Rieffel, 2004)
-For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space.
-Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$.
-We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
-Appendix E Experimental Details and Additional Results
-We provide experimental details below. All the experiments in this paper were executed with Python 3.9 on an Intel(R) Core(TM) i9-10885H 2.4GHz processor with the Windows 10 operating system.
-Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task.
-(Average value of results of five different runs.
-The error bar represents the standard deviation.)
-E.1 Experiment with synthetic data
-We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$.
-We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.
-For $x^{i}$, $y^{i}$, $\tilde{x}^{i}$, and $a$, we discretized each function with 30 equally spaced points on $\mathbb{T}$.
-In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$ discussed in Remark 15, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0.
-Figure 3 shows the result.
-The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$.
-Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
-We also show additional results regarding the deep approach in Figure 4.
-The result is the average $\pm$ the standard deviation of three independent runs.
-We set different values of $n$ with the same deep setting considered in Subsection 7.1.
-Figure 4: Test error of the regression task with deep approach with different $n$ and $L$.
-E.2 Experiment with MNIST
-We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.
-Figure 5 (a) shows the test error.
-We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$.
-In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0.
-Figure 5 (b) shows the result.
-The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$.
-Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
-Figure 5: (a) Test error of the image recovering task with different values of $n$ and the kernel $k=\hat{k}_{n}^{\operatorname{prod},q}$.
-(Box plot of results of five independent runs with different training and test data.) (b) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the image recovering task.
-(Average value of results of five different runs.
-The error bar represents the standard deviation.)
+Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition.